### v1

# Offline Fictitious Self-Play for Competitive Games

Abstract
Offline Reinforcement Learning (RL) has received significant interest due to its ability to improve policies in previously collected datasets without online interactions. Despite its success in the single-agent setting, offline multi-agent RL remains a challenge, especially in competitive games. Firstly, unaware of the game structure, it is impossible to interact with the opponents and conduct a major learning paradigm, self-play, for competitive games. Secondly, real-world datasets cannot cover all the state and action space in the game, resulting in barriers to identifying Nash equilibrium (NE). To address these issues, this paper introduces Off-FSP, the first practical model-free offline RL algorithm for competitive games. We start by simulating interactions with various opponents by adjusting the weights of the fixed dataset with importance sampling. This technique allows us to learn best responses to different opponents and employ the Offline Self-Play learning framework. In this framework, we further implement Fictitious Self-Play (FSP) to approximate NE. In partially covered real-world datasets, our methods show the potential to approach NE by incorporating any single-agent offline RL method. Experimental results in Leduc Hold’em Poker show that our method significantly improves performances compared with state-of-the-art baselines.
1 Introduction
Multi-agent reinforcement learning (MARL) provides a powerful learning framework to tackle the problems in multi-agent systems and has been applied to a wide range of domains, such as Go (Silver et al., 2017), strategy games (Vinyals et al., 2019), robotics (Yu et al., 2023), unmanned aerial vehicle (Yun et al., 2022) and network routings (Ye et al., 2015). MARL requires extensive interaction and exploration with accurate and efficient environments, however, this requirement is often not fulfilled in practice, such as football (Kurach et al., 2020), negotiation (Yang et al., 2020), wildlife security protection (Fang et al., 2016), and crowdsourcing (Gerstgrasser et al., 2021). The existing data-driven method, supervised learning, also called behavior cloning, heavily relies on high-quality datasets, which are commonly expensive and suboptimal. Offline MARL offers an excellent alternative to solve these issues by improving policies from previously collected datasets without further interactions (Tseng et al., 2022; Yang et al., 2021).
Recent research in offline MARL has predominantly concentrated on cooperative tasks, where all agents receive a common reward. For cooperative tasks, the goal can be seen as maximizing the cumulative rewards of a joint policy of all agents. Thus, existing single-agent offline RL methods can be easily extended to multi-agent scenarios  (Tseng et al., 2022). However, competitive tasks, often framed as zero-sum games, necessitate a divergent goal orientation.
In the realm of competitive games, or called zero-sum games, exemplified by Texas hold’em poker in studies by (Brown & Sandholm, 2018, 2019), the pursuit of a Nash equilibrium emerges as a prevalent solution concept. Complementary efforts in online setting by (Silver et al., 2017; Ye et al., 2020) focus on the ELO ratio or returns. To reach these goals, These approaches predominantly rely on the self-play paradigm, wherein policies are continuously refined to maximize returns against evolving adversary policies. Nonetheless, in the offline scenario, self-play doesn’t work due to the absence of online interactions, complicating the development of effective offline learning algorithms for competitive games.
In addressing competitive games, recent benchmarks such as AlphaStar Unplugged (Mathieu et al., 2023) and Hokoff (Qu et al., 2023; Wei et al., 2022) have applied single-agent offline RL algorithms to multi-agent settings. However, maximizing the cumulative rewards against static opponents results in an overfitting policy and vulnerable exploitation by more dynamic opponents. Moreover, when datasets are collected by poor policies, such as random policies, it is pointless to learn to defeat these weak opponents, because real-life opponents typically exhibit rational and high-performance behaviors. Li et al. (2022) proposes a model-based offline paradigm for equilibrium finding but requires a strong assumption that datasets fully cover the state-action space of original games. Cui & Du (2022a); Zhong et al. (2022); Cui & Du (2022b) contribute to theoretical insights for a weaker assumption on datasets but only propose algorithms that are feasible in theory. These works encounter challenges in integration with single-agent offline RL methods and do not offer a practical algorithm for addressing real-world problems.
In this paper, we proposed an offline learning framework, called Offline Self-Play (Off-SP), and an offline learning algorithm, Offline Fictitious Self-Play (Off-FSP), for equilibrium finding in zero-sum extensive-form games, bridging the gap between single-agent offline RL and competitive games. To our knowledge, Off-FSP is the first model-free offline algorithm for practical zero-sum game that offers the flexibility to combine with various Offline RL agents. We first propose a technique to approximate interaction with different opponents by re-weighting the datasets with importance sampling. This allows us to learn the best responses against arbitrary opponents with offline RL and derives self-play under the offline setting. By implementing the fictitious self-play on Off-SP, Off-FSP acquires the ability to learn strategies that approximate NE using fixed datasets. We also discuss the challenges of missing data for real-world datasets. The experiments in an imperfect-information two-player zero-sum game, Leduc Hold’em Poker, show that our method outperforms the state-of-the-art baselines and can approach a Nash equilibrium in datasets that fully cover the state-action space of the game. Even with partially covered datasets, our methods show the ability to reduce exploitability in original games.
2 Related Work
In competitive games, also known as zero-sum games, people have developed a series of learning algorithms, where finding Nash equilibrium with self-play is the major learning paradigm for this problem. Fictitious self-play (FSP) (Heinrich et al., 2015) combines fictitious play (FP) (Brown, 1951) with self-play and provably converges to an NE in extensive-form games. Counterfactual regret minimization (CFR) (Bowling et al., 2015) combines regret minimization with self-play, first solving an imperfect-information game of real-world scale. DeepNash (Perolat et al., 2022) extends regularised Nash dynamics with RL and converges to an approximate NE in Stratego. Policy-Space Response Oracles (PSRO) (Lanctot et al., 2017) learn to find NE by iteratively learning best responses to its policies’ population, which can also be seen as a population-based self-play. Some works (Silver et al., 2017; Ye et al., 2020; Yang et al., 2022) of large-scale games aim for a higher winning percentage or returns and try to exploit opponents, not pursuing strict NE, still based on the self-play paradigm.
In the offline learning paradigm, behavior cloning (BC) is the simplest form that directly imitates the sampling strategy of datasets. The performance of BC is limited by the sampling strategy in the dataset, while offline RL can surpass this limitation by maximizing returns. Offline RL faces the challenge of the extrapolation error on out-of-distribution states and actions. Batch-constrained deep q-learning (BCQ) (Fujimoto et al., 2019b) and conservative q-learning (CQL) (Kumar et al., 2020) mitigate this issue by constraining the gap between policy and data distribution. Critic regularized regression (CRR) (Wang et al., 2020) solve this problem by weighted behavior cloning.
In offline competitive scenarios, BC is a viable algorithm, still limited by the quality of sampling strategy in datasets. Mathieu et al. (2023); Qu et al. (2023) directly learn strategies by single-agent offline RL without self-play. This way also requires high-quality behavior strategies in datasets and is easily overfitted to fixed opponents. Cui & Du (2022a); Zhong et al. (2022); Cui & Du (2022b) propose theoretically feasible algorithms. Li et al. (2022) proposed a model-based learning framework, OEF, for extending online equilibrium-finding algorithms in offline scenarios. However, OEF only supports datasets with full coverage of state-action spaces, which is not realistic for real-world problems. Although previous theoretic methods allow for partially covered datasets, they cannot be practically applied to real-life problems. In this paper, Off-SP enables the self-play paradigm in offline MARL of competitive games. Off-FSP implement a practical algorithm to find NE and has the potential to learn on partially covered datasets by integrating any variants of single-agent offline RL algorithms.
3 Preliminaries
3.1 Extensive-form Game
Extensive-form games are a model of sequential interaction involving $n$ agents. Each player’s goal is to maximize his payoff in the game. At each step $t$ of extensive-form game, only one player observes his respective information states $s^{i}_{t}\in\mathcal{S}^{i}$ and suggests his action $a^{i}_{t}\in\mathcal{A}^{i}(s^{i}_{t})$. $\mathcal{S}^{i}$ is the set of information states of player $i$, and $\mathcal{A}^{i}(s)$ is the set of aviliable action at state $s$. The game’s randomization is defined by an extra player $c$, called chance. Chance follows a fixed randomized strategy that determines the distribution of the next state at its turns. The player function $P:\mathcal{S}\rightarrow\mathcal{N}\cup\{c\}$, with $\mathcal{N}=\{1,\dots,n\}$ denotes the set of players, determines who is to act at a given state.
We assume the games with perfect recall. Each player’s current information state $s_{t}^{i}$ implies the trajectory sequence $(s_{1}^{i},a_{1}^{i},s_{2}^{i},a_{2}^{i},\dots,s_{t}^{i})$, including information states and actions led to $s_{t}^{i}$. In extensive-form games, each player plays following a behavioural strategy that maps all information states to distributions of available actions. The realization-plan (Von Stengel, 1996), $x(s_{t}^{i})=\Pi_{j=1}^{t-1}\pi^{i}(a_{j}^{i}|s_{j}^{i})$, describes the probability of reaching the information state $s_{t}^{i}$ following player $i$’s behavioural strategy, $\pi^{i}$. The strategy profile $\pi=(\pi^{1},\dots,\pi^{n})$, is a joint of all player’s behavioural strategy. $\pi^{-i}$ denotes the strategy profile of all players excepts $i$. The payoff of player $i$ is denoted by $R^{i}(\pi^{i},\pi^{-i})\in\mathbb{R}$, and $\sum_{i}R^{i}=0$ for zero-sum games. Given a fixed $\pi^{-i}$, best response $\textsc{BR}(\pi^{-i})$ is the strategy that has the highest payoff. A Nash equilibrium (NE) is a strategy profile that any strategy $\pi^{i}$ in this profile is a best response to the opponent’s profile $\pi^{-i}$. The $\epsilon$-best response ($\epsilon$-BR) and $\epsilon$-Nash equlibirium ($\epsilon$-NE) are approximations to above defination. $\epsilon$-BR is suboptimal by no more than $\epsilon$ compared with BR. Similarly, $\epsilon$-NE is a profile of $\epsilon$-BR. NashConv (Timbers et al., 2022), also called exploitability in two-player zero-sum games, evaluates the distance from $\pi$ to an NE, defined as $\sum_{i}R^{i}(\textsc{BR}(\pi^{-i}),\pi^{-i})-R^{i}(\pi^{i},\pi^{-i})$.
Normal-form game is a simplified game description in game theory. Normal form represents the game by way of a matrix. An extensive-form game induces an equivalent normal-form game. In Sections 4.4 and 5.1, we discuss with toy examples of a normal-form game, rock-paper-scissors, and show the experimental results.
Algorithm 1 Fictitious Self-Play
Function FictitiousSelfPlay()
      Initialize strategy $\pi_{0}$.  for $k=1$ to $K$ do
           $\mathcal{D}_{k}\leftarrow$ GenerateData($\pi_{k-1}$) foreach player $i\in\mathcal{N}$ do
                $\beta^{i}_{k}\leftarrow$ LearnBestResponse($\mathcal{D}_{k},\pi^{-i}_{k-1}$)
          $\pi_{k}\leftarrow$ UpdateAverageStrategy($\pi^{i}_{k-1},\beta^{1}_{k},\dots,\beta^{n}_{k}$)
     return average strategy $\pi_{K}$
3.2 Fictitious Self-Play
Fictitious Play (Brown, 1951) is a game-theoretic model for learning from self-play. Fictitious players iteratively compute the best responses to opponents’ average behavioural strategy and update their set of strategies. The average strategies of fictitious players are proven to converge to Nash equilibria in two-player zero-sum and many-player potential games. Generalised weakened fictitious play (Leslie & Collins, 2006) is an improvement of fictitious play that enables approximate best responses and perturbed average strategy updates. Generalised weakened fictitious play also provably converges to Nash equilibria in certain game classes and is suitable for machine learning framework.
Original fictitious Play is defined in normal-form games, which requires exponential computations for extensive-form games. Heinrich et al. (2015) introduce Fictitious Self-Play (FSP) that extend fictitious play and generalised weakened fictitious play into the extensive-form game with linear time and space complexity. We briefly describe FSP at  Algorithm 1. In extensive-form games, the average strategies $\pi^{i}_{k}$ are updated by the realization-equivalence theorem. At the $k$-th iteration of FSP, the average behavioural strategy of player $i$ is
$\begin{split}\forall s,a:~{}&\pi^{i}_{k}(a|s)=(1-\lambda)\pi^{i}_{k-1}(a|s)+\lambda\beta^{i}_{k}(a|s),\\ &\lambda=\frac{\alpha_{k}x_{\beta^{i}_{k}}(s)}{(1-\alpha_{k})x_{\pi^{i}_{k-1}}(s)+\alpha_{k}x_{\beta^{i}_{k}}(s)},\end{split}$ (1)
where $\beta^{i}_{k}\in\epsilon_{k}\textsc{-BR}(\pi_{k-1}^{-i})$ is a best-response to opponent $\pi_{k-1}^{-i}$ and $\alpha_{k}$ is parameter to mix the best-response $\beta^{i}_{k}$ and previous average strategy $\pi^{i}_{k-1}$. Strategy $\pi^{i}_{k}$ is equivalent to choosing either strategy $\pi^{i}_{k-1}$ or $\beta^{i}_{k}$ before the beginning of each game, with probabilities $1-\alpha_{k}$ and $\alpha_{k}$ respectively. When $\alpha_{k}\rightarrow 0$ and $\epsilon_{k}\rightarrow 0$ as $k\rightarrow 0$, and $\sum_{k=1}^{\infty}\alpha_{t}=\infty$ for all $n$ players, the average strategy profile $\pi_{k}=(\pi^{1}_{k},\dots,\pi^{n}_{k})$ converge to an NE. In a common setting with $\alpha_{k}=\frac{1}{k}$, $\pi_{K}^{i}$ is equivalent to choosing all strategy $\beta^{i}_{1},\dots,\beta^{i}_{K}$ with the probability of $\frac{1}{K}$. The update of average behavioural strategy requires a traverse for all information state $s^{i}\in\mathcal{S}^{i}$, which is computationally expensive for large games and machine learning-based frameworks. FSP restored all the data of strategies $\beta^{i}_{1},\dots,\beta^{i}_{k}$ and trained an approximate average strategy with supervised learning.
3.3 Offline Reinforcement Learning
(a) Off-FSP and Off-SP.
(b) Differences between FSP and Off-FSP.
Figure 1: Illustration of Off-FSP, Off-SP and three essential functions. The green box in (a) is the Off-SP.
Reinforcement learning (RL) agents aim to maximize the expected cumulative discounted reward in a Markov decision process (MDP). MDP is denoted as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},\rho_{0},r,\gamma,T)$. $\mathcal{S},\mathcal{A}$ represent the state and action spaces. $\mathcal{T}(s_{t+1}|s_{t},a_{t}),r(s_{t},a_{t})$ represent the dynamics and reward function. $\rho_{0}$ is the distribution of initial state $s_{1}\sim\rho_{0}$. $\gamma\in[0,1]$ is the discount factor. Many reinforcement learning algorithms learn from transition tuples, $(s_{t},a_{t},r_{t},s_{t+1})$, where $s_{t},a_{t},r_{t}$ are the state, action and reward at time $t$, and $s_{t+1}$ is the next state after agent take action $a_{t}$ at state $s_{t}$. The expected cumulative discounted reward following policy $\pi$ can be formalized as the action-value function $Q(s_{t},a_{t})=\mathbb{E}_{\pi}[\sum^{\infty}_{i=t}\gamma^{i-t}r(s_{i},a_{i})]$. Model-based RL methods learn the dynamic models while model-free methods directly learn the Q function to state-action pairs. Given the fixed opponent $\pi_{-i}$ and an extensive form game, the game of player $i$ can be defined as a MDP $\mathcal{M}(\pi_{-i})$ (Silver & Veness, 2010; Greenwald et al., 2013). Player $i$’s information state defines the state, and his payoff function defines the reward. An $\epsilon$-optimal policy of the MDP, $\mathcal{M}(\pi^{-i})$, is also the $\epsilon$-BR to the strategy $\pi^{-i}$.
Offline RL algorithm breaks the assumption that the agent can interact with the environment. The offline RL algorithm learns to maximize the expected cumulative reward based on a fixed dataset. The dataset is sampled following a behavior policy $\pi_{b}(a|s)$ in MDP $\mathcal{M}$. $\hat{\pi}_{b}(a|s):=\frac{\sum_{s,a\in\mathcal{D}^{i}}\mathbbm{1}[s=s,a=a]}{\sum_{s\in\mathcal{D}^{i}}\mathbbm{1}[s=s]}$ denote the empirical behavior policy, at all state $s\in\mathcal{D}^{i}$. In offline RL, the Q-function is trained only on actions sampled from offline datasets, which cannot cover all state and action spaces. Therefore, the Q-function may be erroneously overestimated at out-of-distribution (OOD) actions, which is called extrapolation error (Fujimoto et al., 2019b). Recent offline RL methods encourage the policy to learn on the support of training data or employ weighted behavior cloning to mitigate this error.
4 Offline Fictitious Self-Play
In this section, we introduce the Offline Self-Play (Off-SP) learning framework and give an implementation as Offline Fictitious Self-Play (Off-FSP) algorithm to find an $\epsilon$-Nash equilibrium with a fixed dataset. In the online scenario, self-play is a process in which a strategy iteratively maximizes cumulative rewards against itself, and fictitious self-play plays against the average of its past strategies. Corresponding to the concept of the online setting, Off-SP is a learning framework that enables strategy play against changing opponents with a fixed dataset. Off-FSP implement fictitious self-play on this learning framework with average strategies as the changing opponents and learn to find NE.
Initially, we describe the offline datasets and make strong assumptions to simplify the problem. Based on the original FSP, as described in Algorithm 1, we derive the offline fictitious self-play by replacing three essential functions. In Section 4.2, we replace GenerateData by simulating play against different opponents with weighted datasets and replace LearnBestResponse with single-agent offline RL algorithm. The first two functions introduce Off-SP. In Section 4.3, we replace UpdateAverageStrategy by computing the average strategy based on samples and derive Off-FSP. Lastly, we discuss how offline RL helps Off-FSP to reduce the exploitability under real-world datasets.
4.1 Assumptions of Offline Datasets
Trajectory and Dataset.
The trajectory of extensive-form games is a sequence of information states, actions, and player’s id. The trajectory also involves states and actions of chance $c$. The trajectory $\tau_{E}=\{s_{1},a_{1},p_{1},s_{2},\dots,s_{T},a_{T},p_{T}\mid p_{t}=P(s_{t})\},$
We can assume the next information state for a prefix of trajectory and the payoff for a whole trajectory are definite. In other words, if we know the information sets of all players and chance, the game is deterministic.
In player $i$’s perspective, the extensive-form game can be modeled as an MDP $\mathcal{M}$ given a fixed opponent $\pi^{-i}$ (Silver & Veness, 2010; Greenwald et al., 2013). In MDP $\mathcal{M}$, the dynamic function $\mathcal{T}(s^{i}_{t+1}\mid s^{i}_{t},a^{i}_{t})$ models the dynamics of opponents and the chance player. The observed MDP is non-deterministic because of imperfect information in a single player’s information state.
To learn best-response with single-agent RL, we project the trajectory $\tau_{E}$ into player $i$’s perspective with a projection function $\tau^{i}=\mathcal{F}^{i}(\tau_{E})$. The projection $\mathcal{F}^{i}$ filters out the states corresponding to player $i$ while relabeling the time indices.
$\begin{split}\tau^{i}&=\mathcal{F}^{i}(\tau_{E})=\{s_{t},a_{t}\mid\forall p_{t}\in\tau_{E},p_{t}=i\}\\ &=\{s_{1}^{i},a_{1}^{i},s_{2}^{i},\dots,s_{T^{\prime}}^{i},a_{T^{\prime}}^{i}\},\end{split}$
where $T^{\prime}$ is the length of the player $i$’s trajectory $\tau^{i}_{\mathcal{M}}$. The subscripts of $s_{t}$ and $s_{t}^{i}$ are different and represent the time indices in the corresponding trajectories. For state $s^{i}_{t}\in\tau^{i}_{\mathcal{M}}$, we use function $I(s^{i}_{t})$ to denote the subscript of state in $\tau_{E}$. We denote opponent $j$’s latest state at player $i$’s state $s_{t}^{i}$ with function $\tau_{<}^{j}(s_{t}^{i})$.
$\tau_{<}^{j}(s_{t}^{i})=s_{k},~{}\text{where}~{}{k=\max\{k^{\prime}<I(s^{i}_{t})\mid p_{k^{\prime}}=j\}}.$
In order to offer a clear explanation of the stated components, we present an example trajectory in Figure 2.
Figure 2: An example of extensive-form games and datasets. (Left) An example trajectory in extensive form game; (Right) Illustration of trajectories and dataset in extensive-form game and single-agent perspective. The orange area represents the player function. Blue parts represent the trajectories. The purple area represents the projection relationship under function $\mathcal{F}^{1}$. The green parts explain how function $\tau^{2}_{<}$ works.
Based on the description of trajectory in both extensive-form games and single-agent perspective, the datasets consist of multiple corresponding trajectories. For extensive-form game, the dataset is $\mathcal{D}_{E}=\{\tau_{E}\}$. The dataset for player $i$ is $\mathcal{D}^{i}=\{\tau^{i}_{\mathcal{M}}=\mathcal{F}^{i}(\tau_{E})\mid\tau_{E}\in\mathcal{D}_{E}\}$. The $i$-th player’s dataset, derived from dataset $\mathcal{D}_{E}$, is denoted as $\mathcal{D}^{i}=\mathcal{F}^{i}(\mathcal{D}_{E})$. For briefly, we ignore the reward transformation in trajectories.
Assumptions for Offline Learning.
To introduce our method, we first assume the offline dataset $\mathcal{D}_{E}$ is both a fully covered dataset and a real-equivalence dataset. This assumption, crucial for the theoretical derivation of our methods, is not a prerequisite for practical application. The adaptability of our method to real-world datasets, unconstrained by these assumptions, is discussed in Section 4.4. The definitions of these two concepts are provided as follows.
Definition 4.1 (Fully Covered Dataset).
[…]
When a dataset is fully covered, the trained learning algorithm will not suffer from OOD problems.
Definition 4.2 (Real-Equivalence Dataset).
For a dataset of extensive-form game $\mathcal{D}_{E}$ sampled following a behavior strategy $\pi_{b}$, dataset $\mathcal{D}_{E}$ is an real-equivalence dataset if $Pr(\tau_{E})=\frac{\sum_{\tau\in\mathcal{D}_{E}}\mathbbm{1}[\tau=\tau_{E}]}{|\mathcal{D}_{E}|},\forall\tau_{E}\in\mathcal{D}_{E},$ where $Pr(\tau_{E})$ is the probability of sampling trajectory $\tau_{E}$ with $\pi_{b}$ online.
When a dataset is a real-equivalence dataset , sampling trajectories from the offline dataset $\mathcal{D}_{E}$ is equivalent to sampling from the extensive-form game with strategy $\pi_{b}$ online and the right-hand side of the equation is the probability density of $\tau_{E}$ in $\mathcal{D}_{E}$, denoted by $\mathcal{D}_{E}(\tau_{E})$.
4.2 Learning Open-ended Best-Response
Self-play and fictitious self-play iteratively interact with the game to generate data and utilize the data to learn best responses. In $k$-th iteration, the opponent for player $i$ is changed to $\pi^{-i}_{k-1}$, therefore, in order to learn the best response $\beta^{i}_{k}=\text{BR}(\pi^{-i}_{k-1})$, player $i$ must interact with $\pi^{-i}_{k-1}$ to generate new data $\mathcal{D}^{i}_{k}$. The process, from the perspective of player $i$, is equivalent to interacting with a new MDP $\mathcal{M}(\pi^{-i}_{k})$ and maximizing returns with RL to learn the corresponding best response (Greenwald et al., 2013).
In offline settings, we are limited to using a fixed dataset sampled by behavioural policy $\pi_{b}$, in which RL methods can only obtain the best response $\text{BR}(\pi_{b}^{-i})$ for player $i$. In order to execute the fictitious self-play, we should generate player $i$ datasets under different MDP $\mathcal{M}(\pi^{-i}_{k})$. With importance sampling (Nachum et al., 2019), we can emulate sampling from another dataset $\mathcal{D}^{i}_{w}$ with weighting $w(d_{t})=\frac{\mathcal{D}^{i}_{w}(d_{t})}{\mathcal{D}^{i}(d_{t})}$ (Hong et al., 2023) on the original dataset $\mathcal{D}^{i}$, where $\mathcal{D}^{i}_{w}(d_{t})$ and $\mathcal{D}^{i}(d_{t})$ denote the probability density of tuple $d_{t}=(s^{i}_{t},a^{i}_{t},s^{i}_{t+1})$ in according datasets. This formulation gives the following equivalence:
$\mathbb{E}_{d_{t}\sim\mathcal{D}^{i}_{w}}[L(d_{t};\theta)]\Leftrightarrow\mathbb{E}_{d_{t}\sim\mathcal{D}^{i}}[w(d_{t})L(d_{t};\theta)],$ (2)
where $L(d_{t};\theta)$ is the loss function of an off-policy RL method, and $\theta$ is the parameter of RL policy to be optimized. With the assumption that dataset $\mathcal{D}_{E}$ is both a fully covered dataset and a real-equivalence dataset, sampling from it with importance sampling is equivalent to sampling from the online game with importance sampling.
Theorem 4.3.
For player $i$, the weight of transferring opponent from $\pi^{-i}_{b}$ to $\pi^{-i}$ is:
$w(d_{t})=\frac{x^{-i}_{\pi^{-i}}(s_{j}){\pi^{-i}}(a_{j}|s_{j})}{x^{-i}_{\pi_{b}^{-i}}(s_{j}){\pi_{b}^{-i}}(a_{j}|s_{j})},~{}s_{j}=\tau_{<}^{-i}(s_{t+1}^{i}).$ (3)
In order to simplify the phrase, we denote the strategies of $n-1$ opponents as an ensemble strategy $\pi^{-i}$ and refer to them as player $-i$. The proof of Theorem 4.3 is provided in Appendix A. Given an offline dataset $\mathcal{D}_{E}$, the behavior strategy $\pi_{b}$ can be approximated by an empirical behavioural policy $\tilde{\pi}_{b}$. In extensive-form games, we can estimate $\tilde{\pi}_{b}$ by counting or supervised learning policy.
When considering a particular opponent, the weight $w(d_{t})$ assigned to a sample $d_{t}$ may shift to zero or a large value. As a result, when training RL with batched samples, the RL loss suffers from a large variance (Munos et al., 2016), leading to unstable training. To address this issue, we employ $\frac{w(d_{t})}{\sum_{d\in\mathcal{D}}w(d)}$ as the sampling probability rather than multiply the loss function by $w(d_{t})$ as shown in Equation 2. In this way, re-sampling from fully covered real-equivalence datasets is identical to sample data in online games. With batches of data sampled from the dataset, we can apply an offline RL to learn best responses.
Figure 0(b) and Figure 3 illustrates these two functions and Off-SP. In the function GenerateDataset, we first calculate $w(d)$ for all the samples $d_{t}$ and get a re-weighted dataset $\mathcal{D}_{k}^{i}$ for each player $i$. In the function LearnBestResponse, offline RL algorithms repeatedly sample a batch of data and optimize the learned policy for $M$ times. These two functions derive the Off-SP. At each step $t$, Off-SP first generates dataset play against the current opponents and updates strategies towards the best response of the opponents.
Figure 3: Offline Self-Play learning framework.
4.3 Aggregate Average Strategy in Samples
Before diving into Off-FSP, it is essential to review the mechanics of fictitious play in an online scenario. In norm-form games, fictitious play updates the new average strategy $\pi_{k+1}$ following Equation 1 after learning the best responses. To extend FP on extensive-form games, FSP gathers the trajectories of $\beta^{i}_{k}$ during each iteration and employs supervised learning to approximate $\pi^{i}_{k+1}$ for each player. However, Off-FSP does not require the current average strategy $\pi_{k}$ to sample new data, but only requires probability $\pi_{k}(s,a)$ for all state-action pairs in a fixed dataset. Therefore, an interactable average strategy $\pi_{k}$ is not necessary. We can just maintain the probability $\pi_{k}(s,a)$ in the datasets for the weighting technique. Traverse along the trajectory $\tau_{i}$, Off-FSP computes the probability $\pi^{i}_{k}(s,a)$ for player $i$ following Equation 1. Figure 0(b) shows one step of the function UpdateAverageStrategy. To facilitate the calculation, we save the probability into the current dataset $\mathcal{D}^{i}_{k}$, changing the sample into $d^{\prime}=(s,a,\pi^{i}_{k}(s,a))$.
Although the weighting technique does not require a real instance of average behavioural strategy, Off-FSP requires an interactable strategy as output. One way is utilizing supervised learning to imitate the behavioural strategy with the weighting technique. When the dataset is not a fully covered dataset, this restricts strategies to exclusively choosing actions that are present in the dataset. In our experiments, we opted for an alternative method that involves aggregating all policies within a collection $\Pi$. Before the beginning of each game, one of the strategies in $\Pi$ is chosen with a specific probability and the payoff of the aggregation is the expectative payoff over all possible strategies. This approach maintains the option of picking actions that are out of the distribution but necessitates a larger storage capacity for storing all strategies. Algorithm 2 and Figure 0(a) present the pipeline of Offline Fictitious Self-play.
Algorithm 2 Offline Fictitious Self-Play
Function Off-FictitiousSelfPlay($\mathcal{D}_{E}$):
      Estimate the empirical behavioural policy $\tilde{\pi}_{b}$$\Pi_{i}\leftarrow\{\pi_{b}^{i}\},\quad\pi_{0}\leftarrow\tilde{\pi}_{b}$$\mathcal{D}^{i}_{0}\leftarrow\{(s,a,\pi^{i}_{0}(s,a))|(s,a)\in\mathcal{F}^{i}(\mathcal{D}_{E})\}$for $k=1$ to $K$ do
           for each player $i\in\mathcal{N}$ do
                $\mathcal{D}^{i}_{k}\leftarrow$ GenerateDataset($\mathcal{D}_{E},\mathcal{D}^{-i}_{k-1}$) $\beta^{i}_{k}\leftarrow$ LearnBestResponse($\mathcal{D}^{i}_{k},\beta^{i}_{k-1}$) $\mathcal{D}^{i}_{k}\leftarrow$ UpdateAverageStrategy($\mathcal{D}^{i}_{k-1},\beta_{k}^{i}$) Update collection of BRs $\Pi_{i}\leftarrow\Pi_{i}\cup\{\beta_{k}^{i}\}$
     return Aggregate($\Pi$)
Function GenerateDataset($\mathcal{D}_{E},\mathcal{D}^{-i}_{k-1}$):
[…]
Function LearnBestResponse($\mathcal{D}^{i}_{k},\beta^{i}_{k-1}$):
      Initialize policy parameters with $\beta^{i}_{k}\leftarrow\beta^{i}_{k-1}$  Optimize $\beta^{i}_{k}$ with offline RL and data $\mathcal{D}^{i}_{k}$return $\beta^{i}_{k}$
Function UpdateAverageStrategy($\mathcal{D}^{i}_{k-1},\beta^{i}_{k}$):
      for $(s,a,\pi_{k-1}^{i}(s,a))\in\mathcal{D}^{i}_{k-1}$ do
           Calculate $\pi_{k}^{i}(s,a)$ with $\beta^{i}_{k},\pi^{i}_{k-1}$ following Equation 1, and update dataset $\mathcal{D}^{i}_{k}$
     return $\mathcal{D}^{i}_{k}$
4.4 Learning in Real-world Datasets
Considering an offline dataset sampled from extensive-form games, the two assumptions in Section 4.1 cannot be completely fulfilled. Initially, a discrepancy in the probability of trajectories is invariably present between game and sampled datasets. This results in an irremovable error in learning best responses. Leslie & Collins (2006) proposed weakened fictitious play that allows suboptimal $\epsilon$-best responses in the learning process. The discrepancy does not totally make fictitious play fail but affects the distance of the final solution from equilibrium. In practice, sampling a large dataset can mitigate the issue. The sample complexity is a common problem discussed in previous works about offline learning (Cui & Du, 2022a; Yan et al., 2022).
Secondly, it is not realistic to gather a fully covered dataset, especially in large-scale games. Previous works have discussed the requirements of the coverage to find NE (Cui & Du, 2022a; Li et al., 2022). However, their assumption is too restrictive to be fulfilled in real-world problems. In contrast, this work discusses how to minimize the distance to NE, i.e. the exploitability, under a given dataset, which is a more realistic setting. In fictitious self-play, a straightforward way to reduce exploitability is to learn better $\epsilon$-best responses at each step, which is the same for Off-FSP. The LearnBestResponse function can utilize any variants of the offline RL algorithms, which effectively address the issue of incomplete data set coverage.
Figure 4: Example partially covered datasets for Rock-Paper-Scissors. Numbers in the grids show the probability density of different samples. The red dashed boxes indicate the probability of different actions for corresponding behavioural strategies. (Left) A partially covered example with infinite samples. (Right) A partially covered example with finite samples.
Considering infinite samples, the missing of some $(s,a)$ is caused by $\pi_{b}(s,a)=0$ where $\pi_{b}$ is the behavior strategy to sample the dataset. We show this case in a simple matrix game, Rock-Paper-Scissors(RPS), with the first example of Figure 4. In this case, the covered part of the matrix can be seen as a subgame, where agents can only choose actions from observed ones. The payoff of the subgame is known. A conservative solution for this dataset is to find the Nash equilibrium of the subgame. However, if the lower bound of payoff in this game is known, allowing the selection of out-of-distribution actions may be a more reasonable setting. With the knowledge that payoffs of RPS are in the range of $[-1,1]$, player $1$ consistently selects paper as it yields a payoff of $1$. For player $2$, the only action, rock, in the dataset gets a payoff of $-1$, and choosing OOD actions may be a better selection because the rock is already the worst action. Some offline RL algorithms, such as CQL, give the OOD actions low rewards rather than set hard restrictions on these actions, which enables the learned policy to choose OOD actions. The experimental results in  Sections 5.1 and 5.2 show that restricting the policy space to subgames leads to higher exploitability.
[…]
5 Experiments
In this section, we design a series of experiments to show the performance of Off-FSP and investigate the OOD issue in offline competitive games. We first evaluate Off-FSP on toy examples in a matrix game, rock-paper-scissors. We compared Off-FSP against the state-of-the-art offline RL methods for competitive games, including OEF (Li et al., 2022), Behavior Cloning (BC), and other single-agent Offline RL methods, on an imperfect-information poker game, Leduc Hold’em. We also conduct ablation studies on Off-FSP in  Section D.2.
To demonstrate the compatibility of our approach with existing single-agent offline RL algorithms, we combine Off-FSP with three methods: CQL (Kumar et al., 2020), BCQ (Fujimoto et al., 2019b) and CRR (Wang et al., 2020). For variants with CQL, denoted as CQL-$\alpha$, we conduct different hyperparameters $\alpha$ which indicates the penalty level of choosing OOD actions. Higher $\alpha$ means less OOD actions performed by the policy. When $\alpha=0$, there is no penalty, and CQL degrades to DQN. More details about the experiments are provided in Appendix C.
5.1 Rock, Paper, Scissors
To provide empirical evidence for our statements in Section 4.4 and show the effectiveness of our methods, we first give results on three datasets of rock-paper-scissors. The first dataset, FULL, is randomly sampled from a uniform behavioural strategy, $\pi_{b}=(\frac{1}{3},\frac{1}{3},\frac{1}{3})$, which is a fully covered dataset and an approximation of ideal dataset. P1 and P2 are partially covered datasets described in Figure 4. Figure 5 shows the NashConv within $500$ iterations and the average strategies of player $2$ after the end of training.
In FULL dataset, both CQL-00 and CQL-$1$ with Off-FSP are converged to a $\epsilon$-NE with a small gap with online fictitious play. In the partially covered P2 dataset, only CQL with $\alpha>0$ can properly address the OOD issues successfully converges to a $\epsilon$-NE, while CQL-00 is misleading by one single sample of Rock2 and selects Rock2 with a probability of $1$. In the P1 scenario, employing CQL with a smaller $\alpha$ value results in a reduced NashConv. This outcome is attributed to the allowance of OOD actions, as elaborated in Section 4.4. The outcomes from experiments P1 and P2 demonstrate the complexities inherent in OOD problems. The findings suggest that neither arbitrary selection nor total exclusion of OOD actions is optimal, emphasizing the need for sophisticated offline algorithms.
Figure 5: Results of RPS. (Left) NashConv. (Right) Player $2$’s Average Strategy.
5.2 Leduc Hold’em
(a) NashConv
(b) OOD ratio
(c) NashConv w/o OOD actions
(d) Payoff
Figure 6: Results of Leduc Hold’em
We now show the results compared with existing algorithms for offline competitive games. We collect 5 datasets on Leduc Hold’em with different quality. Mix and Random are nearly fully covered datasets, which are sampled by a random policy and aggregate strategy with expert and random policy. P20, P10, and P5 are partially covered datasets sampled from mixed strategies of $20,10,5$ best responses. These best responses are randomly picked from the training process of online PSRO. The more BR picked for sampling, the more state-action pairs are covered by the datasets. We visualize the ratio of covered state-action pair for all datasets in Appendix D. All these datasets contain 10 0001000010\,00010 000 trajectories. Mix and P20 datasets sampled from a high-quality strategy, so we can easily obtain a good strategy by supervised learning their empirical behavioral strategies $\tilde{\pi}_{b}$. This paradigm is also called behavior cloning, which is a popular baseline for offline learning.
We evaluate state-of-the-art offline learning baselines and variants of Off-FSP on all the datasets. Figure 5(a) shows the NashConv of all the algorithms, and all variants of our method outperform other baselines. The result also demonstrates the ability of Off-FSP to robustly integrate with any existing offline RL methods. BC supervised learning the empirical behavioral strategies $\tilde{\pi}_{b}$ of all datasets, which shows Off-FSP can also outperform the sampling strategy. Mix and P20 datasets are both sampled from high-quality strategies, while the Random dataset is sampled from a random strategy. OEF-DCFR and OEF-PSRO are two variants of OEF (Li et al., 2022), which is a model-based offline algorithm for finding NE. OEF mixed its strategy with BC, evaluated it in online games multiple times and used the minimum NashConv as its results. This operation involves interaction with online games, which violates the offline scenario. To fair comparison, we remove the mixing operation in our main experiment. The results with mixing are shown in  Appendix D. OEF operates based on the assumption of full data coverage, so OEF strategies have a high probability of performing OOD actions on partially covered datasets, as shown in Figure 5(b).
To investigate the influence of OOD actions, we restrict all the strategies on the observed state-action pairs and evaluate these restricted strategies. Figure 5(c) demonstrates significant performance degradation in some methods. We also learned an online CFR to find a $\epsilon$-NE on the subgames of different datasets. For partially covered datasets, the $\epsilon$-NE shows large NashConv values. This indicates that directly solving NE in subgames is insufficient for resolving the offline learning problem in competitive games.
In previous works (Mathieu et al., 2023; Qu et al., 2023), directly applying single-agent Offline RL on the fixed dataset is also an attempt to solve offline competitive games. This paradigm is identical to learning best responses to the behavioural strategies $\pi_{b}$. These strategies overfit their opponents and are vulnerable to exploitation. These works only pay attention to strategies with higher returns and do not discuss NE. Therefore, we show the payoff of variants of Off-FSP playing with corresponding single-agent Offline RL algorithms, as Figure 5(d). Our methods show positive payoff to all the single-agent methods.
6 Conclusion
In this paper, we study offline multi-agent reinforcement learning for competitive games and propose Off-SP and Off-FSP to enable single-agent offline RL algorithms to be applied in this scenario. We find that Off-FSP can approximate NE even with partially covered datasets. Extensive experiments show that all variants of Off-FSP significantly outperform state-of-the-art baselines in different datasets of an imperfect-information two-player zero-sum game, Leduc Hold’em.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Derivation of Theorem
Proof.
Now we derive the weight when changing the opponent to $\pi_{-i}$. Since we do not need to adjust player $i$’s strategy, player $i$’s strategy is still considered as $\tilde{\pi}_{b}^{i}$. In player $i$’s perspective, the reach probability of a tuple $d_{t}=(s^{i}_{t},a^{i}_{t},s^{i}_{t+1})$ is
$Pr(d_{t}|\pi^{-i})=\Pi_{j=0}^{t}\pi^{i}(s^{i}_{j},a^{i}_{j})\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})\cdot\pi^{i}(a^{i}_{t+1}|s^{i}_{t+1}),$
where $\mathcal{T}_{\pi^{-i}}$ is the transition function under MDP $\mathcal{M}(\pi^{-i})$. So we have:
$w_{\pi^{-i}}(d_{t})$ $=\frac{Pr(d_{t}|\pi^{-i})}{Pr(d_{t}|\tilde{\pi}^{-i}_{b})}$
$=\frac{\Pi_{j=0}^{t}\tilde{\pi}_{b}^{i}(s^{i}_{j},a^{i}_{j})\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})}{\Pi_{j=0}^{t}\tilde{\pi}_{b}^{i}(s^{i}_{j},a^{i}_{j})\mathcal{T}_{\tilde{\pi}_{b}^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})}\cdot\frac{\tilde{\pi}_{b}^{i}(a^{i}_{t+1}|s^{i}_{t+1})}{\tilde{\pi}_{b}^{i}(a^{i}_{t+1}|s^{i}_{t+1})}$
$=\Pi_{j=0}^{t}\frac{\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})}{\mathcal{T}_{\tilde{\pi}_{b}^{-i}}(s^{i}_{j+1}|s_{j},a_{j})}$
Considering $\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})$ and $\mathcal{T}_{\tilde{\pi}_{b}^{-i}}(s^{i}_{j+1}|s_{j},a_{j})$, both of these two items are the multiplication of a series of transition probabilities, of which only the probability of opponents’ strategies $\pi^{-i}$ is different. So the above equation can be simplified to:
$w_{\pi^{-i}}(d_{t})$ $=\Pi_{j=0}^{t}\frac{{\pi^{-i}}(a_{j}|s_{j})}{{\tilde{\pi}_{b}^{-i}}(a_{j}|s_{j})}$
$=\frac{x_{\pi^{-i}}(s_{j}){\pi^{-i}}(a_{j}|s_{j})}{x_{\tilde{\pi}_{b}^{-i}}(s_{j}){\tilde{\pi}_{b}^{-i}}(a_{j}|s_{j})},$
where $s_{j}=\tau_{<}^{-i}(s_{t+1}^{i})$. ∎
Appendix B Baselines
Behaviour Cloning (BC).
Behaviour Cloning directly learns the probability distribution of actions conditioned on the observation from the dataset, hoping to recover the performance of the behavior policy used to generate the dataset.
Conservative Q-Learning (CQL)
Conservative Q-Learning (Kumar et al., 2020) learns a conservative Q-function such that the expected value of a policy under this Q-function lower-bounds its actual value in order to alleviate overestimation of values induced by the distributional shift between dataset and the learned policy. In practice, we solve such an optimization problem to update our Q-function, where the first term is the penalty for lower-bounding the Q-function, and the second term is the Bellman error. We use CQL based on the QRDQN.
$\min_{Q}\alpha\mathbb{E}_{\mathbf{s}\sim\mathcal{D}}\left[\log\sum_{\mathbf{a}}\exp(Q(\mathbf{s},\mathbf{a}))-\mathbb{E}_{\mathbf{a}\sim\hat{\pi}_{\beta}(\mathbf{a}\mid\mathbf{s})}[Q(\mathbf{s},\mathbf{a})]\right]+\frac{1}{2}\mathbb{E}_{\mathbf{s},\mathbf{a},\mathbf{s}^{\prime}\sim\mathcal{D}}\left[\left(Q-\hat{\mathcal{B}}^{\pi_{k}}\hat{Q}^{k}\right)^{2}\right]$ (4)
Critic Regularized Regression (CRR)
Critic Regularized Regression (Wang et al., 2020) handles the problem of offline policy optimization by value-filtered regression. It selectively imitates the dataset by choosing a function $f$ that is monotonically increasing in $Q_{\theta}$.
[…]
Batch-Constrained deep Q-learning (BCQ)
Batch-Constrained deep Q-learning (Fujimoto et al., 2019b) uses a generative model to learn the distribution of transitions in the dataset, samples several actions from it in the current state, perturbates them and chooses the one with highest Q value. In this way, they tried to minimize the distance of selected actions to the dataset and lead to states similar to those in the dataset. In the experiments, we use a discrete BCQ variant introduced in (Fujimoto et al., 2019a).
Single-Agent Offline RL
In this set of baseline methods, each player independently learns a policy from the dataset using single-agent offline RL algorithms, including CQL, CRR, and BCQ. It’s theoretically equivalent to figuring out the best response to the fixed opponent policy used to collect the dataset. This kind of method is also employed in Qu et al. (2023) and Mathieu et al. (2023).
Offline Equilibrium Finding (OEF)
Offline Equilibrium Finding (Li et al., 2022) is a model-based framework to find the game equilibrium in the offline setting. They directly train a model to describe the game dynamics and apply online equilibrium-finding algorithms like PSRO and DCFR (a deep learning version of CFR) to compute equilibrium. They also combine the behavior cloning policy with the model-based policy for improvement, where they mix these two policies with different weights and evaluate each weight in the online game to find the best combination.
Appendix C Details of experiments
C.1 Datasets
C.1.1 Rock, Paper, Scissors
FULL is randomly sampled from a uniform behavioral strategy, $\pi_{b}=(\frac{1}{3},\frac{1}{3},\frac{1}{3})$, i.e., every player randomly picks each action with the same probability. P1 and P2 are generated according to the exact proportion described in Figure 4, e.g., P1 comprises 500 episodes of Rock to Rock and 500 episodes of Paper to Rock.
Every dataset is composed of 1000 trajectories.
C.1.2 Leduc Hold’em
Random is sampled by uniformly randomly picking action from all legal actions in every decision node.
Mix is sampled from a mixed strategy, which is a 3-to-1 mix of expert and random strategies. The expert strategy is trained by PSRO with 30 iterations and is a mix of 30 deterministic policies according to a specific ratio.
Partial dataset (5-10-20) is sampled by first picking 5, 10, and 20 policies from the 30 components of the expert strategy and then uniformly mixing them to form the sampling strategy.
Every dataset comprises 10000 trajectories except those used in ablation study about the dataset size.
C.2 Variants of Off-FSP
Off-FSP-CQl, Off-FSP-CRR, and Off-FSP-CQL are different variants of our methods that integrate with different single-agent offline RL methods in the function of LearnBestResponse. We then introduce details about these Offline RL methods.
In CQL, we use a QRDQN as the base algorithm and a simple MLP as the Q network. Hence, the network’s input is the observation vector, and the output has a dimension of the number of quantiles times the number of actions. The hidden layer of MLP contains the same number of neurons. We use ReLU as the activation function between every linear layer.
In CRR, the actor and critic networks have the same structure, first a shared feature network and then a MLP head. The shared feature network is a MLP with two layers, and the size of the second layer is the number of actions. All activation functions are ReLU. CQL penalty term is also added to the CRR objective function for better performance.
In BCQ, the policy network and the imitation network have the same structure, which is also first a shared feature network and then an MLP head, exactly the same as CRR.
C.3 Hyperparameters
In OEF, all settings are the same as Li et al. (2022).
In RPS, we use Adam optimizer with a learning rate of 0.005 and betas (0.9, 0.999), and the network has only one layer, and the number of quantiles in QRDQN is 100. The target network is not used. In each iteration, the network is trained five epochs and is updated 100 times each epoch with a batch size of 1000.
In Leduc, the hyperparamters are shown in Table 1. All networks are optimized by Adam Optimizer (Kingma & Ba, 2017). The hidden layer number of CRR and BCQ describes the whole network, including the feature network and MLP head. The target network is used.
Table 1: Hyperparameters used in Leduc
Hyperparameter CQL CRR BCQ
Hidden layer number 4 5 6
Hidden layer size 256 128 256
Learning rate 0.0001 0.0001 0.0001
Adam $(\beta_{1},\beta_{2})$ (0.9, 0.999) (0.9, 0.999) (0.9, 0.999)
Training steps per iteration $M$ 1 000 1 000 1 000
Target Network update frequency 100 100 100
Batch size 1024 1024 1024
QRDQN number of quantiles 100
CQL $\alpha$ 0.1 0.1
BCQ unlikely action threshold 0.1
BCQ imitation logits penalty 0.01
CRR improve mode Exponential
CRR $\beta$ 1
CRR ratio bound 20
Appendix D Supplement Results
D.1 Analysis of Datasets
Figure 7: State-Action Cover Ratio
We visualize the state-action cover ratio of different datasets in Figure 7
D.2 Ablations
(a) Value of $\alpha$
(b) Size of Data
Figure 8: Results of ablations.
[…]
D.3 Learning Curves of Off-FSP
OEF evaluates strategies’ performance by allowing strategies to mix with BC by different weights. However, the results are decided by the best-mixed strategies. This is equivalent to allowing strategies to be interactively evaluated online, which is contrary to the setting of offline scenarios. And we cannot tell whether this result is caused by the original strategy or BC. We still provide the final results under this evaluation method in Figure 9, and our method still outperforms OEF.
Figure 9: NashConv with evaluation method in OEF

### v2

@@
 Abstract
-Offline Reinforcement Learning (RL) has received significant interest due to its ability to improve policies in previously collected datasets without online interactions.
+Offline Reinforcement Learning (RL) enables policy improvement from fixed datasets without online interactions, making it highly suitable for real-world applications lacking efficient simulators.
 Despite its success in the single-agent setting, offline multi-agent RL remains a challenge, especially in competitive games.
@@
 We start by simulating interactions with various opponents by adjusting the weights of the fixed dataset with importance sampling.
-This technique allows us to learn best responses to different opponents and employ the Offline Self-Play learning framework.
-In this framework, we further implement Fictitious Self-Play (FSP) to approximate NE.
-In partially covered real-world datasets, our methods show the potential to approach NE by incorporating any single-agent offline RL method.
-Experimental results in Leduc Hold’em Poker show that our method significantly improves performances compared with state-of-the-art baselines.
+This technique allows us to learn the best responses to different opponents and employ the Offline Self-Play learning framework.
+To overcome the challenge of partial coverage, we combine the single-agent offline RL method with Fictitious Self-Play (FSP) to approximate NE by constraining the approximate best responses away from out-of-distribution actions.
+Experiments on matrix games, extensive-form poker, and board games demonstrate that Off-FSP achieves significantly lower exploitability than state-of-the-art baselines.
+Finally, we validate Off-FSP on a real-world human-robot competitive task, demonstrating its potential for solving complex, hard-to-simulate real-world problems.
 1 Introduction
-Multi-agent reinforcement learning (MARL) provides a powerful learning framework to tackle the problems in multi-agent systems and has been applied to a wide range of domains, such as Go (Silver et al., 2017), strategy games (Vinyals et al., 2019), robotics (Yu et al., 2023), unmanned aerial vehicle (Yun et al., 2022) and network routings (Ye et al., 2015).
-MARL requires extensive interaction and exploration with accurate and efficient environments, however, this requirement is often not fulfilled in practice, such as football (Kurach et al., 2020), negotiation (Yang et al., 2020), wildlife security protection (Fang et al., 2016), and crowdsourcing (Gerstgrasser et al., 2021).
-The existing data-driven method, supervised learning, also called behavior cloning, heavily relies on high-quality datasets, which are commonly expensive and suboptimal.
-Offline MARL offers an excellent alternative to solve these issues by improving policies from previously collected datasets without further interactions (Tseng et al., 2022; Yang et al., 2021).
-Recent research in offline MARL has predominantly concentrated on cooperative tasks, where all agents receive a common reward.
-For cooperative tasks, the goal can be seen as maximizing the cumulative rewards of a joint policy of all agents.
-Thus, existing single-agent offline RL methods can be easily extended to multi-agent scenarios  (Tseng et al., 2022).
-However, competitive tasks, often framed as zero-sum games, necessitate a divergent goal orientation.
-In the realm of competitive games, or called zero-sum games, exemplified by Texas hold’em poker in studies by (Brown & Sandholm, 2018, 2019), the pursuit of a Nash equilibrium emerges as a prevalent solution concept.
-Complementary efforts in online setting by (Silver et al., 2017; Ye et al., 2020) focus on the ELO ratio or returns.
-To reach these goals, These approaches predominantly rely on the self-play paradigm, wherein policies are continuously refined to maximize returns against evolving adversary policies.
-Nonetheless, in the offline scenario, self-play doesn’t work due to the absence of online interactions, complicating the development of effective offline learning algorithms for competitive games.
-In addressing competitive games, recent benchmarks such as AlphaStar Unplugged (Mathieu et al., 2023) and Hokoff (Qu et al., 2023; Wei et al., 2022) have applied single-agent offline RL algorithms to multi-agent settings.
+Multi-agent reinforcement learning (MARL) provides a powerful learning framework to tackle the problems in multi-agent systems and has been applied to a wide range of domains, such as Go (Silver et al.
+2017), strategy games (Vinyals et al.
+2019), robotics (Yu et al.
+2023), unmanned aerial vehicle (Yun et al.
+2022), and network routings (Ye, Zhang, and Yang 2015).
+MARL typically relies on extensive interaction and exploration with accurate and efficient environments, hindering its application in the real world.
+In many real-world multi-agent problems, such as football (Kurach et al.
+2020), negotiation (Yang, Chen, and Narasimhan 2020), and crowdsourcing (Gerstgrasser, Trivedi, and Parkes 2021), the absence of reliable simulators makes it expensive and inefficient to interact with the environment and collect new data for training MARL agents.
+Moreover, in special problems, such as wildlife protection (Fang et al.
+2016), learning MARL agents online by trial and error is unsafe, which could lead to danger for patrols or wildlife.
+In these cases, learning from existing data can be extremely valuable as it does not require additional sampling.
+Offline MARL offers an excellent alternative to solve these issues by improving policies from previously collected datasets without further interactions (Tseng et al.
+2022; Yang et al.
+2021).
+Figure 1: Comparison of offline self-play with other learning paradigms for real-world competitive tasks.
+Online self-play requires a simulator and suffers from the sim-to-real gap.
+Behavior cloning needs costly expert data.
+Naive offline RL optimizes with incorrect objectives.
+In contrast, Off-SP learns from inexpensive, low-quality data.
+See Section˜5.2 for experiments on the illustrated real-world game.
+In competitive multi-agent games, often framed as zero-sum games, the objective is to find a Nash equilibrium (NE) (Kreps 2018) or maximizing the ELO ratio (Silver et al.
+2017; Ye et al.
+2020), which is divergent from the goal of maximizing cumulative rewards in single-agent situations.
+Exemplified by studies of Texas hold’em poker (Brown and Sandholm 2018, 2019), the pursuit of a Nash equilibrium enables agents to learn robust policies against various opponents, enhancing their performance in competitive environments.
+To achieve the goals, online approaches predominantly rely on the self-play paradigm (Zhang et al.
+2024), wherein policies are continuously refined to maximize returns against evolving adversary policies.
+However, in the offline scenario, the challenge arises from the absence of online interactions with evolving opponents, complicating the development of self-play paradigms.
+Recent benchmarks of offline competitive games, such as AlphaStar Unplugged (Mathieu et al.
+2023) and Hokoff (Qu et al.
+2023; Wei et al.
+2022), have directly applied single-agent offline RL algorithms with no self-play.
 However, maximizing the cumulative rewards against static opponents results in an overfitting policy and vulnerable exploitation by more dynamic opponents.
 Moreover, when datasets are collected by poor policies, such as random policies, it is pointless to learn to defeat these weak opponents, because real-life opponents typically exhibit rational and high-performance behaviors.
+The reliance on high-quality or high-coverage datasets, which are commonly expensive and suboptimal in the real world, is another challenge for offline learning.
+The existing data-driven method, supervised learning, also called behavior cloning, ignores the objective of online learning and only imitates the sampling policy of datasets.
+Consequently, this method performs poorly when dealing with non-expert datasets.
+Many recent works are learning to improve policies toward the NE, but they demand high state-action coverage within the dataset.
 Li et al.
-(2022) proposes a model-based offline paradigm for equilibrium finding but requires a strong assumption that datasets fully cover the state-action space of original games.
-Cui & Du (2022a); Zhong et al.
-(2022); Cui & Du (2022b) contribute to theoretical insights for a weaker assumption on datasets but only propose algorithms that are feasible in theory.
-These works encounter challenges in integration with single-agent offline RL methods and do not offer a practical algorithm for addressing real-world problems.
-In this paper, we proposed an offline learning framework, called Offline Self-Play (Off-SP), and an offline learning algorithm, Offline Fictitious Self-Play (Off-FSP), for equilibrium finding in zero-sum extensive-form games, bridging the gap between single-agent offline RL and competitive games.
-To our knowledge, Off-FSP is the first model-free offline algorithm for practical zero-sum game that offers the flexibility to combine with various Offline RL agents.
+(2022) proposes a model-based offline paradigm for equilibrium finding, but it requires a strong assumption that datasets fully cover the state-action space of the original games.
+Cui and Du (2022b); Zhong et al.
+(2022); Cui and Du (2022a) contribute to theoretical insights for a weaker assumption on datasets but only propose theoretically feasible algorithms.
+Existing work lacks a practical method applicable to non-expert and partially covered real-world datasets.
+Figure˜1 compares our paradigm with existing works, highlighting the advantage of Off-SP on solving real-world competitive problems.
+Detailed comparisons with related works are provided in the Appendix˜B.
+In this paper, we propose an offline learning framework, called Offline Self-Play (Off-SP), and an offline learning algorithm, Offline Fictitious Self-Play (Off-FSP), for equilibrium finding in zero-sum extensive-form games, bridging the gap between single-agent offline RL and competitive games.
+To our knowledge, Off-FSP is the first model-free offline algorithm for practical zero-sum games that offers the flexibility to combine with various Offline RL agents and improve policies on non-expert real-world datasets.
 We first propose a technique to approximate interaction with different opponents by re-weighting the datasets with importance sampling.
-This allows us to learn the best responses against arbitrary opponents with offline RL and derives self-play under the offline setting.
-By implementing the fictitious self-play on Off-SP, Off-FSP acquires the ability to learn strategies that approximate NE using fixed datasets.
-We also discuss the challenges of missing data for real-world datasets.
-The experiments in an imperfect-information two-player zero-sum game, Leduc Hold’em Poker, show that our method outperforms the state-of-the-art baselines and can approach a Nash equilibrium in datasets that fully cover the state-action space of the game.
-Even with partially covered datasets, our methods show the ability to reduce exploitability in original games.
-2 Related Work
+This allows us to learn the approximate best responses against arbitrary opponents with offline RL and derive a self-play paradigm under the offline setting.
+Also, for partially covered and non-expert datasets, we use a surrogate loss, NashConv, to measure the distance to NE, rather than finding the exact NE.
+Off-FSP combines single-agent offline reinforcement learning methods with fictitious self-play to learn approximate best responses iteratively and minimize the NashConv.
+We validate our approach across matrix-form and extensive-form zero-sum games, demonstrating consistently low exploitability and superior performance over existing offline RL baselines under partially covered datasets.
+Notably, we further apply our method to a real-world human-robot competitive task, showcasing its potential in addressing hard-to-simulate decision-making problems.
+2 Preliminaries
+2.1 Extensive-form Game
+Extensive-form games are a model of sequential interaction involving $n$ agents.
+Each player’s goal is to maximize his payoff in the game.
+At each step $t$ of extensive-form game, only one player observes his respective information states $s^{i}_{t}\in\mathcal{S}^{i}$ and suggests his action $a^{i}_{t}\in\mathcal{A}^{i}(s^{i}_{t})$.
+$\mathcal{S}^{i}$ is the set of information states of player $i$, and $\mathcal{A}^{i}(s)$ is the set of available action at state $s$.
+The player function $\mathcal{P}:\mathcal{S}\rightarrow\mathcal{N}$, with $\mathcal{N}=\{1,\dots,n\}$ denotes the set of players, determines the player to act.
+In extensive-form games, each player plays following a policy $\pi^{i}:\mathcal{S}^{i}\rightarrow\Delta(\mathcal{A}^{i})$ that maps information states to distributions of actions.
+The realization-plan (Von Stengel 1996), $x(s_{t}^{i})=\Pi_{j=1}^{t-1}\pi^{i}(a_{j}^{i}|s_{j}^{i})$, describes the probability of reaching the information state $s_{t}^{i}$ following player $i$’s policy, $\pi^{i}$.
+The strategy profile $\pi=(\pi^{1},\dots,\pi^{n})$, is a joint of all player’s policy.
+$\pi^{-i}$ denotes the strategy profile of all players excepts $i$.
+The payoff of player $i$ is denoted by $R^{i}(\pi^{i},\pi^{-i})\in\mathbb{R}$, and $\sum_{i}R^{i}=0$ for zero-sum games.
+Given a fixed $\pi^{-i}$, best response $\textsc{BR}(\pi^{-i})$ is the policy with the highest payoff.
+A Nash equilibrium (NE) is a strategy profile that any policy $\pi^{i}$ in this profile is a BR to the opponent’s profile $\pi^{-i}$.
+The $\epsilon$-best response ($\epsilon$-BR) and $\epsilon$-Nash equlibirium ($\epsilon$-NE) are approximations to the above definition.
+$\epsilon$-BR is suboptimal by no more than $\epsilon$ compared with BR.
+Similarly, $\epsilon$-NE is a profile of $\epsilon$-BR.
+NashConv (Timbers et al.
+2022), also called exploitability in two-player zero-sum games, evaluates the distance from $\pi$ to an NE, defined as $\sum_{i}R^{i}(\textsc{BR}(\pi^{-i}),\pi^{-i})-R^{i}(\pi^{i},\pi^{-i})$.
+2.2 Fictitious Self-Play
+Fictitious Self-Play (FSP)(Heinrich, Lanctot, and Silver 2015) is a game-theoretic model that iteratively computes the best responses to opponents’ average policy and updates their set of policies.
+We briefly describe FSP at  Appendix˜A.
+In extensive-form games, the average policies $\pi^{i}_{k}$ are updated by the realization-equivalence theorem.
+At the $k$-th iteration of FSP, the average policy $\pi^{i}_{k}$ of player $i$ is
+$\begin{split}\forall s,a:~&\pi^{i}_{k}(a|s)=(1-\lambda)\pi^{i}_{k-1}(a|s)+\lambda\beta^{i}_{k}(a|s),\\ \lambda&=\frac{\alpha_{k}x_{\beta^{i}_{k}}(s)}{(1-\alpha_{k})x_{\pi^{i}_{k-1}}(s)+\alpha_{k}x_{\beta^{i}_{k}}(s)},\end{split}$ (1)
+where $\beta^{i}_{k}\in\epsilon_{k}\textsc{-BR}(\pi_{k-1}^{-i})$ is a best-response to opponent $\pi_{k-1}^{-i}$ and $\alpha_{k}$ is the mixing parameter.
+Policy $\pi^{i}_{k}$ is equivalent to choosing either policy $\pi^{i}_{k-1}$ or $\beta^{i}_{k}$ before the beginning of each game, with probabilities $1-\alpha_{k}$ and $\alpha_{k}$ respectively.
+A standard choice is $\alpha_{k}=\frac{1}{k}$.
+Figure 2: Example Datasets of RPS.
+Numbers in the grids show the probability density of different samples.
+The red dashed boxes indicate the probability of different actions for corresponding behavioural policies.
+Figure 3: Results of RPS. The prefix of D1- and D2- are refering to restuls on the first and second datasets respectively.
+2.3 Offline Reinforcement Learning
+Reinforcement learning (RL) agents aim to maximize the expected cumulative discounted reward in a Markov decision process (MDP).
+MDP is denoted as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},\rho_{0},r,\gamma,T)$.
+$\mathcal{S},\mathcal{A}$ represent the state and action spaces.
+$\mathcal{T}(s_{t+1}|s_{t},a_{t}),r(s_{t},a_{t})$ represent the dynamics and reward function.
+$\rho_{0}$ is the distribution of initial state $s_{1}\sim\rho_{0}$.
+$\gamma\in[0,1]$ is the discount factor.
+The expected cumulative discounted reward following policy $\pi$ can be formalized as the action-value function $Q(s_{t},a_{t})=\mathbb{E}_{\pi}[\sum^{\infty}_{i=t}\gamma^{i-t}r(s_{i},a_{i})]$.
+Given the fixed opponent $\pi_{-i}$ and an extensive form game, the game of player $i$ can be defined as a MDP $\mathcal{M}(\pi_{-i})$ (Silver and Veness 2010; Greenwald et al.
+2013).
+An $\epsilon$-optimal policy of the MDP, $\mathcal{M}(\pi^{-i})$, is also the $\epsilon$-BR to the policy $\pi^{-i}$.
+Offline RL algorithm breaks the assumption that the agent can interact with the environment.
+The offline RL algorithm learns to maximize the expected cumulative reward based on a fixed dataset.
+The dataset is sampled following a behavior policy $\pi_{b}(a|s)$ in MDP $\mathcal{M}$.
+$\hat{\pi}_{b}(a|s):=\frac{\sum_{s,a\in\mathcal{D}^{i}}\mathds{1}[s=s,a=a]}{\sum_{s\in\mathcal{D}^{i}}\mathds{1}[s=s]}$ denote the empirical behavior policy, at all state $s\in\mathcal{D}^{i}$.
+The Q-function may be erroneously overestimated at out-of-distribution (O.
+O.
+D.) actions, which is called extrapolation error (Fujimoto, Meger, and Precup 2019).
+Recent offline RL methods encourage the policy to learn on the support of training data or employ weighted behavior cloning to mitigate this error.
+3 The Motivating Example: Rock-Paper-Scissors
+We start by analysing the well-known game, Rock-Paper-Scissors (RPS), to illustrate the challenges of learning in offline datasets of competitive games.
+In this game, two players can choose one of three actions: Rock, Paper, or Scissors.
+The payoffs are defined as follows: Rock beats Scissors, Scissors beats Paper, and Paper beats Rock.
+The game has a unique Nash equilibrium (NE) where both players play each action with equal probability of $\frac{1}{3}$.
+In Figure˜3, we show two datasets of RPS.
+The first dataset is a fully covered dataset sampled from a non-expert policy, $P(\text{R,P,S})=(0.6,0.2,0.2)$.
+The second dataset is a partially covered dataset sampled from a modified asymmetric RPS game, where the second player has a new action, Rock2, with the same payoff as Rock.
+In the first non-expert dataset, neither Behavioral Cloning (BC) nor naive offline RL can find the NE, which highlights the importance of the self-play paradigm.
+BC mimics the distribution of the dataset by supervised learning.
+Its policy, $P_{\text{BC}}(\text{R,P,S})=(0.6,0.2,0.2)$, is suboptimal and beaten by the policy of playing Paper only.
+As the solution of  Mathieu et al.
+(2023); Qu et al.
+(2023), single-agent offline RL maximizes the payoff under the fixed opponent in the dataset, and its policy, $P_{RL}(\text{R,P,S})=(0,1,0)$, is also easy to be exploited by another policies.
+Under the self-play paradigm, the policy learns to play against a dynamic opponent, which is more robust and can approximate the NE.
+Figure˜3 compares the performance of our method, Off-FSP-CQL, with the baselines, BC, Off-FSP-DQN and online Fictitious Play (FP) (Brown 1951).
+Corresponding to the challenge of learning in partially covered datasets, the second dataset in Figure˜3 leaves the payoff of Rock2 not fully observed.
+In such datasets, finding the accurate NE of the original game is impossible , so the goal of learning is to minimize the exploitability and NashConv, i.e., the distance to NE.
+In the dataset, the only observed sample of Rock2 is $(\text{Scissors},\text{Rock2})$.
+In player $2$’s perspective, the Rock2 action gets a payoff of $1$ all the time, so choosing it with a probability of $1$ is the best choice.
+Nevertheless, the policy has the highest exploitability in the original game, as player $1$ possesses a best response by selecting scissors, resulting in a payoff of $1$ for player $1$.
+Offline RL algorithms, such as Conservative Q-Learning(CQL) (Kumar et al.
+2020), mitigate this problem by punishing the agent for learning OOD actions, including unseen actions and undersampled actions.
+By combining self-play and offline RL, Off-FSP ignores the samples with large uncertainty and learns robust policies with low exploitability.
+As shown in  Figure˜3, Off-FSP without offline RL (Off-FSP-DQN) converges to high exploitability, while Off-FSP with CQL can reduce the exploitability to a low level.
+We also show the learning curve of an online algorithm, FP, which shows the convergence speed of Off-FSP is comparable to online algorithms.
+Further explanations of the results are given in the Appendix˜E.
+((a)) Offline Self-Play framework.
+((b)) Pipeline of Off-FSP.
+((c)) Differences between FSP and Off-FSP.
+Figure 4: Illustration of Off-SP, Off-FSP and three essential steps. The green box in (b) is Off-SP.
+4 Offline Fictitious Self-Play
+In this section, we introduce the Offline Self-Play (Off-SP) learning framework and give an implementation as Offline Fictitious Self-Play (Off-FSP) algorithm to minimize NashConv, the distance to NE, with a fixed dataset.
+Off-SP learns policies iteratively, maximizing cumulative rewards against changing opponents without interactable environments.
+Off-FSP adopts the fictitious self-play on Off-SP, where the policy plays against the average of past opponents.
+Initially, we describe the offline datasets and make assumptions to simplify the problem.
+Based on the original FSP, as described in Appendix˜A, we derive the offline fictitious self-play with modifications on three essential functions, GenerateData, LearnBestResponse, and UpdateAveragePolicy.
+In Off-SP, GenerateData simulates play against different opponents with weighted datasets, and LearnBestResponse optimizes the policy with an existing single-agent offline RL algorithm.
+Offline RL algorithm, such as CQL, ensures the performance of BR and reduction of NashConv even on partially covered datasets.
+In Section˜4.3, we introduce UpdateAveragePolicy by computing the average policy on samples and derive Off-FSP.
+4.1 Problem Formulation
+The trajectory of extensive-form games is a sequence of information states, actions, and rewards.
+The trajectory is $\tau_{E}=\{s_{1},a_{1},r_{1},\dots,s_{T},a_{T},r_{T}\}$.
+In player $i$’s perspective, the extensive-form game can be modeled as an MDP $\mathcal{M}$ given a fixed opponent $\pi^{-i}$ (Silver and Veness 2010; Greenwald et al.
+2013).
+In MDP $\mathcal{M}$, the dynamic function $\mathcal{T}(s^{i}_{t+1}\mid s^{i}_{t},a^{i}_{t})$ models the dynamics of opponents.
+The goal of Off-FSP is to iteratively learn the best-response policy $\pi^{i}$ against changing opponents and minimize NashConv.
+To learn best-response with single-agent RL, we project the trajectory $\tau_{E}$ into player $i$’s perspective with a projection function $\tau^{i}=\mathcal{F}^{i}(\tau_{E})$.
+The projection $\mathcal{F}^{i}$ filters out the states corresponding to player $i$ while relabeling the time indices.
+$\begin{split}\tau^{i}=\mathcal{F}^{i}(\tau_{E})&=\{s_{t},a_{t},r_{t}\mid\forall s_{t}\in\tau_{E},\mathcal{P}(s_{t})=i\}\\ &=\{s_{1}^{i},a_{1}^{i},r_{1}^{i},s_{2}^{i},\dots,s_{T^{\prime}}^{i},a_{T^{\prime}}^{i},r_{T^{\prime}}^{i}\},\end{split}$
+where $T^{\prime}$ is the length of the player $i$’s trajectory $\tau^{i}_{\mathcal{M}}$.
+The subscripts of $s_{t}$ and $s_{t}^{i}$ are different and represent the time indices in the corresponding trajectories.
+For state $s^{i}_{t}\in\tau^{i}_{\mathcal{M}}$, we use function $I(s^{i}_{t})$ to denote the subscript of state in $\tau_{E}$, and $\tau_{<}^{j}(s_{t}^{i})$ denotes opponent $j$’s latest state at $s_{t}^{i}$.
+$\tau_{<}^{j}(s_{t}^{i})=s_{k},~\text{where}~{k=\max\{k^{\prime}<I(s^{i}_{t})\mid p_{k^{\prime}}=j\}}.$
+In order to offer a clear explanation of the stated components, we present an example trajectory in Figure 5.
+Based on the description of trajectory in both extensive-form games and single-agent perspective, the datasets consist of multiple corresponding trajectories.
+For extensive-form game, the dataset is $\mathcal{D}_{E}=\{\tau_{E}\}$.
+The dataset for player $i$ is $\mathcal{D}^{i}=\{\tau^{i}_{\mathcal{M}}=\mathcal{F}^{i}(\tau_{E})\mid\tau_{E}\in\mathcal{D}_{E}\}$.
+The $i$-th player’s dataset, derived from dataset $\mathcal{D}_{E}$, is denoted as $\mathcal{D}^{i}=\mathcal{F}^{i}(\mathcal{D}_{E})$.
+Figure 5: An illustration example of trajectory $\tau^{1}$ and $\tau_{E}$.
+Purple parts represent Player 1.
+The yellow arrows imply the projection relationship under function $\mathcal{F}^{1}$.
+Green part indicates $\tau_{<}^{1}$.
+To introduce the theoretical foundation of our method, we begin by defining an idealized dataset condition that facilitates analysis.
+These assumptions are not required for practical application and are used only to isolate and clarify the core mechanisms of our approach.
+In later sections, we show how the method generalizes to realistic datasets with partial coverage by integrating standard offline RL algorithms.
+Definition 4.1 (Fully Covered Dataset).
+$\mathcal{S}\text{~and~}\mathcal{A}$ represent the joint sets of information states and actions for all players $\mathcal{N}$.
+A dataset $\mathcal{D}_{E}$ in extensive-form games is a fully covered dataset if $\forall s\in\mathcal{S},a\in\mathcal{A}(s)\Rightarrow(s,a)\in\mathcal{D}_{E}$.
+When a dataset is fully covered, the trained learning algorithm will not suffer from OOD problems.
+Definition 4.2 (Real-Equivalence Dataset).
+For a dataset of extensive-form game $\mathcal{D}_{E}$ sampled following a policy $\pi_{b}$, dataset $\mathcal{D}_{E}$ is an real-equivalence dataset if $Pr(\tau_{E})=\frac{\sum_{\tau\in\mathcal{D}_{E}}\mathds{1}[\tau=\tau_{E}]}{|\mathcal{D}_{E}|},\forall\tau_{E}\in\mathcal{D}_{E},$ where $Pr(\tau_{E})$ is the probability of sampling trajectory $\tau_{E}$ with $\pi_{b}$.
+When a dataset is a real-equivalence dataset , sampling trajectories from the offline dataset $\mathcal{D}_{E}$ is equivalent to sampling from the extensive-form game with policy $\pi_{b}$ online and the right-hand side of the equation is the probability density of $\tau_{E}$ in $\mathcal{D}_{E}$, denoted by $\mathcal{D}_{E}(\tau_{E})$.
+4.2 Offline Self-Play
+Self-play iteratively interacts with the game to generate data and utilizes the data to learn the best responses.
+In $k$-th iteration, the opponent for player $i$ is changed to $\pi^{-i}_{k-1}$, therefore, in order to learn the best response $\beta^{i}_{k}=\text{BR}(\pi^{-i}_{k-1})$, player $i$ must interact with $\pi^{-i}_{k-1}$ to generate new data $\mathcal{D}^{i}_{k}$.
+The process of learning BR, from the perspective of player $i$, is equivalent to interacting with a new MDP $\mathcal{M}(\pi^{-i}_{k})$ and maximizing returns with RL (Greenwald et al.
+2013).
+In offline settings, we are limited to using a fixed dataset sampled by behavioural policy $\pi_{b}$, in which RL methods can only obtain the best response $\text{BR}(\pi_{b}^{-i})$ for player $i$.
+In order to execute the fictitious self-play, we generate player $i$’s datasets under different MDP $\mathcal{M}(\pi^{-i}_{k})$.
+With importance sampling (Nachum et al.
+2019), we can emulate sampling from another dataset $\mathcal{D}^{i}_{w}$ with weighting $w(d_{t})=\frac{\mathcal{D}^{i}_{w}(d_{t})}{\mathcal{D}^{i}(d_{t})}$ (Hong et al.
+2023) on the original dataset $\mathcal{D}^{i}$, where $\mathcal{D}^{i}_{w}(d_{t})$ and $\mathcal{D}^{i}(d_{t})$ denote the probability density of tuple $d_{t}=(s^{i}_{t},a^{i}_{t},s^{i}_{t+1})$.
+This formulation gives the following equivalence:
+$\mathbb{E}_{d_{t}\sim\mathcal{D}^{i}_{w}}[L(d_{t};\theta)]\Leftrightarrow\mathbb{E}_{d_{t}\sim\mathcal{D}^{i}}[w(d_{t})L(d_{t};\theta)],$ (2)
+where $L(d_{t};\theta)$ is the loss function of an off-policy RL method, and $\theta$ is the parameter of the RL policy to be optimized.
+With the assumption that dataset $\mathcal{D}_{E}$ is both a fully covered dataset and a real-equivalence dataset, sampling from it with importance sampling is equivalent to sampling from the online game with importance sampling.
+Theorem 4.3.
+For player $i$, the weight of transferring the opponent from $\pi^{-i}_{b}$ to $\pi^{-i}$ is:
+$w(d_{t})=\frac{x^{-i}_{\pi^{-i}}(s_{j}){\pi^{-i}}(a_{j}|s_{j})}{x^{-i}_{\pi_{b}^{-i}}(s_{j}){\pi_{b}^{-i}}(a_{j}|s_{j})},~s_{j}=\tau_{<}^{-i}(s_{t+1}^{i}).$ (3)
+The proof of Theorem 4.3 is provided in Appendix C.
+Given an offline dataset $\mathcal{D}_{E}$, the policy $\pi_{b}$ can be approximated by an empirical behavioural policy $\tilde{\pi}_{b}$.
+We can estimate $\tilde{\pi}_{b}$ by counting or supervised learning policy.
+In the learning process, the weight $w(d_{t})$ assigned to a sample $d_{t}$ may shift to zero or a large value.
+As a result, the RL loss suffers from a large variance (Munos et al.
+2016), leading to unstable training.
+To address this issue, we employ $\frac{w(d_{t})}{\sum_{d\in\mathcal{D}}w(d)}$ as the sampling probability rather than multiplying the loss function by $w(d_{t})$ as shown in Equation 2.
+In this way, re-sampling from fully covered real-equivalence datasets is identical to sampling data in online games.
+With batches of data sampled from the dataset, we can apply an offline RL to learn BR.
+Figures˜4(c) and 4(a) illustrates these two functions and Off-SP.
+In the function GenerateData, we first calculate $w(d)$ for all the samples $d_{t}$ and get a re-weighted dataset $\mathcal{D}_{k}^{i}$ for each player $i$.
+In the function LearnBestResponse, offline RL algorithms repeatedly sample a batch of data and optimize the learned policy for $M$ times.
+We use CQL (Kumar et al.
+2020) as the default offline RL algorithm in our implementation, i.e., $\beta^{i}_{k}=\text{CQL}(\mathcal{D}_{k}^{i})$.
+These two functions derive the Off-SP.
+At each step $t$, Off-SP first generates a dataset playing against the current opponents and updates policies towards the best response of the opponents.
+Figure 6: Results on Extensive-Form Games. (Top) NashConv on Mix Datasets; (Bottom) NashConv on Population Datasets.
+Figure 7: Coverage ratio of datasets.
+4.3 Aggregate Average Policy in Samples
+Similar to FSP, Off-FSP play against the average policy $\pi_{k+1}$ following  Equation˜1.
+Although an interactable average policy $\pi_{k}$ is not necessary in the offline setting, we can just maintain the probability $\pi_{k}(s,a)$ in the datasets for the weighting technique.
+Traverse along the trajectory $\tau_{i}$, Off-FSP computes the probability $\pi^{i}_{k}(s,a)$ for player $i$ following Equation˜1.
+Figure˜4(c) shows one step of the function UpdateAveragePolicy.
+To facilitate the calculation, we save the probability into the current dataset $\mathcal{D}^{i}_{k}$, changing the sample into $d^{\prime}=(s,a,\pi^{i}_{k}(s,a))$.
+In short, we have:
+$\pi^{i}_{k}(s,a)\leftarrow\frac{k-1}{k}x^{i}_{\pi^{i}_{k-1}}(s)\pi^{i}_{k-1}(s,a)+\frac{1}{k}x^{i}_{\beta^{i}_{k}}(s)\beta^{i}_{k}(s,a).$
+In the evaluation phase, we aggregate all policies within a collection $\Pi$.
+Before the beginning of each game, one of the policies in $\Pi$ is chosen with a specific probability, and the payoff of the aggregation is the expected payoff over all possible policies.
+To keep the storage efficient, we only keep policies at fixed interval steps for the real-world application in Section˜5.2.
+Algorithm 2 in Appendix˜A and Figure˜4(b) presents the pipeline of Offline Fictitious Self-play.
+Memory and Computational Complexity.
+Compared with offline RL, Off-FSP only assigns additional weights for samples in the dataset, so the memory complexity is still $O(|\mathcal{D}_{E}|)$ in training.
+The training time of Off-FSP can be noted as $T=T_{reweight}+T_{learn}$, where $T_{reweight}$ is the time of re-weighting and $T_{learn}$ is the time of offline RL learning.
+In empirical experiments of Section˜5, the $T_{reweight}:T_{learn}\approx 1:1$.
+5 Experiments
+In this section, we design experiments to show the performance of Off-FSP in offline competitive games.
+Three benchmark extensive-form games are selected for analysis: Leduc Poker, Large Kuhn Poker, and Oshi Zumo, which include two stochastic poker games and one deterministic board game.
+These environments allow efficient sampling and easy NashConv computation, facilitating a comprehensive assessment of algorithm performance.
+We compare Off-FSP with state-of-the-art baselines, including OEF (Li et al.
+2022) and Behavior Cloning (BC).
+We further conduct ablations on offline RL components to demonstrate the flexibility of Off-FSP.
+Finally, the practical applicability of Off-FSP is highlighted through deployment in a real-world human-robot competitive task, showcasing its potential for solving complex, hard-to-simulate problems.
+Further experimental details are provided in the appendix.
+5.1 Extensive-form Games
+To further evaluate Off-FSP on complicated extensive-form games, we collected datasets of different quality on Leduc Poker, Large Kuhn Poker, and Oshi Zumo.
+Following the previous work (Li et al.
+2022), the first type of datasets is called mix datasets, which are sampled from a mixture of expert and random policies.
+Evaluation on random, mixed, and expert datasets is also a common practice in single-agent offline RL (Kumar et al.
+2020).
+We sample 5 different mix datasets for each game, the ratio of sampling from the expert are $0,0.25,0.5,0.75,1$.
+The expert policy is learned by online PSRO (Lanctot et al.
+2017) with 40 iterations.
+The second type of dataset is called the population dataset, which is sampled from the population of online PSRO.
+$5$ population datasets are uniformly sampled from all population policies of $1,10,20,30,40$ iterations of online PSRO.
+This type of dataset is designed to simulate the datasets sampled from a population of people with different levels of expertise and is similar to the setting in the real world.
+The population datasets with $1$ iteration and the mix datasets with a ratio of 0 are random datasets.
+Every dataset comprises 10 000 trajectories, which is far less than the number in the online PSRO.
+We visualize the coverage of terminal states, i.e., leaf nodes, in Figure˜7.
+Most of the datasets are partially covered.
+Figure˜7 shows the NashConv of Off-FSP and baselines on three extensive-form games.
+The difficulty of learning from datasets is affected by two factors: the quality of sampling policies and the coverage of datasets.
+NashConvs of BC show the quality of sampling policies.
+In most cases, Off-FSP shows the best performance, and the performance of Off-FSP is also robust to both the quality of sampling policies and the coverage of datasets.
+Both OEF-DCFR and OEF-PSRO, two variants of OEF (Li et al.
+2022), fail in most cases.
+OEF tries to find an offline equilibrium with the model-based paradigm, but it is easy to be misled by O.O.D.
+actions.
+OEF mixed its policy with BC, evaluated it in online games multiple times, and used the minimum NashConv as a result.
+To make a fair comparison, we removed the mixing operation in OEF from our main experiment.
+The results with mixing are shown in  Appendix˜G.
+Compared with BC, Off-FSP outperforms in non-expert datasets and achieves comparable performance in expert datasets.
+In expert datasets, the diversity of samples is limited, and few samples can be exploited to improve the policy.
+Off-FSP aims to improve the policy with non-expert datasets, which is more practical and easier to scale up in real-world problems.
+Figure 8: Results of ablation study on random and mix data of Leduc Poker.
+Ablation Studies.
+In previous experiments, we showed the performance of Off-FSP with CQL as the offline RL algorithm to learn the best response.
+However, the choice of an offline RL algorithm is also an important factor.
+Figure˜8 shows learning curve of Off-FSP with different RL algorithms, including CQL, BCQ (Fujimoto, Meger, and Precup 2019), CRR (Wang et al.
+2020), and DQN on two datasets of Leduc Poker.
+Without the constraint of offline RL to avoid O.O.D.
+actions, Off-FSP with DQN converges to a policy with high exploitability.
+Corresponding to the design of Off-FSP, it has the potential to combine with any offline RL algorithms.
+Off-FSP with CQL, BCQ, and CRR show similar performances in these two datasets.
+5.2 Real-world Human-Robot Confrontation
+Figure 9: Setups of the human-robot competitive game. G1 to G4 in sub-figure (d) are four actions of the human player.
+To further evaluate Off-FSP, we design a complex real-world zero-sum game—a human-robot ball defence task, which serves as a simplified version of football defense and attack.
+Figures˜1 and 9 illustrate the task scenario.
+This setup involves a human player, making it difficult to simulate and costly to collect data, thus highlighting the significance of Off-FSP.
+The game consists of two players: a human thrower, who attempts to throw a ball into one of three baskets (goals), and a robotic arm acting as a goalkeeper that tries to block the ball.
+The human thrower receives a reward of +1+1 if the ball enters a goal and 1-1 otherwise.
+As a zero-sum game, the robot receives the negative of the human’s reward and uses a black racket to defend the goals.
+The robot is equipped with a stereo RGBD camera that estimates the ball’s position in real time (see Figure˜9c).
+Its observation space includes the ball’s position and velocity, as well as the position of the robot’s end effector.
+The robot policy outputs target end-effector positions at $60$ Hz.
+These are executed via inverse kinematics and a PD controller.
+For the human player, we simplify their action space to four discrete targets: the three baskets (G1 to G3) or outside the basket area (G4), as shown in Figure˜9d.
+The target is solely determined by the ball’s flight trajectory after release and does not change whether the robot blocks it.
+Therefore, G4 corresponds to imprecise human throws and represents a poor strategy.
+The setup of this task is shown in Figure˜9.
+We collected a mixed dataset comprising $50$ expert trajectories and $150$ low-quality trajectories.
+Expert data were collected with a human operator teleoperating the robot arm while another expert performed the throws.
+The average win rate of the teleoperated robot in these expert trajectories was $64\%$.
+Low-quality trajectories were collected by deploying the robot with a random policy, which moves to uniformly sampled positions within a constrained action space.
+The human thrower was not an expert and occasionally missed the targets (G4).
+This protocol makes low-quality data cheaper to collect, requiring only a non-expert human thrower.
+Methods G1 G2 G3 Worst Average
+BC 0.55 0.35 0.4 0.35 0.43
+BC-Expert 0.45 0.5 0.65 0.45 0.53
+CQL 0.4 0.85 0.8 0.4 0.68
+Off-FSP-CQL 0.8 0.75 0.9 0.75 0.82
+Table 1: Winning rate of robot goalkeeper in ball defense game. The colored numbers represent the worst case between goals.
+Since one player is human, we evaluate only the robot’s policy across different algorithms.
+As NashConv is difficult to compute in real-world settings, we adopt the robot’s win rate as the evaluation metric.
+We compare Off-FSP-CQL with three baselines: BC, BC-Expert, and CQL.
+Here, BC-Expert denotes BC policy trained only on trajectories with a reward of +1+1.
+OEF failed in this task, so we removed it from the comparison.
+For each goal, a total of 20 throws were performed collectively by 10 human throwers (2 throws per thrower on average).
+As shown in Table˜1, Off-FSP-CQL outperforms all baselines with consistently robust performance.
+The dataset shows an imbalanced distribution $P(\text{G1,G2,G3,G4})=(0.29,0.11,0.32,0.28)$, with relatively fewer throws to the middle goal (G2) and a high proportion of off-target throws (G4).
+Off-FSP learns a more balanced human policy $(0.33,0.22,0.40,0.04)$, leading to improved robustness against diverse human strategies.
+By focusing only on high-reward data, BC-Expert and CQL overfit to average opponents in the dataset and fail to defend consistently across all goal targets (G1–G3).
+6 Conclusion
+In this paper, we study offline multi-agent reinforcement learning for competitive games and propose Off-SP and Off-FSP to enable single-agent offline RL algorithms to be applied in this scenario.
+We find that Off-FSP can approximate NE even with partially covered datasets.
+Extensive experiments show that all variants of Off-FSP significantly outperform state-of-the-art baselines in different datasets of multiple two-player zero-sum games.
+Appendix A Algorithms
+Algorithm˜1 describes Fictitious Self-Play (FSP) in the online setting.
+Algorithm˜2 is the Offline Fictitious Self-Play.
+Compared with FSP, the function GenerateData and UpdateAveragePolicy only require two additional traverses of the dataset and skip the time of interacting with the environment, so it does not introduce too much computational overhead.
+1
+2
+3Function FictitiousSelfPlay()
+4   Initialize policy $\pi_{0}$.;
+5
+6 for $k=1$ to $K$ do
+7
+8    $\mathcal{D}_{k}\leftarrow$ GenerateData($\pi_{k-1}$)
+9    foreach player $i\in\mathcal{N}$ do
+10       $\beta^{i}_{k}\leftarrow$ LearnBestResponse($\mathcal{D}_{k},\pi^{-i}_{k-1}$)
+11      end foreach
+12    $\pi_{k}\leftarrow$ UpdateAveragePolicy($\pi^{i}_{k-1},\beta^{1}_{k},\dots,\beta^{n}_{k}$)
+13   end for
+14 return average policy $\pi_{K}$;
+15
+16 end
+Algorithm 1 Fictitious Self-Play
+1
+2
+3Function Off-FictitiousSelfPlay($\mathcal{D}_{E}$):
+4   Estimate the empirical behavioural policy $\tilde{\pi}_{b}$;
+5
+6 $\Pi_{i}\leftarrow\{\pi_{b}^{i}\},\quad\pi_{0}\leftarrow\tilde{\pi}_{b}$;
+7
+8 $\mathcal{D}^{i}_{0}\leftarrow\{(s,a,\pi^{i}_{0}(s,a))|(s,a)\in\mathcal{F}^{i}(\mathcal{D}_{E})\}$;
+9
+10 for $k=1$ to $K$ do
+11    for each player $i\in\mathcal{N}$ do
+12       $\mathcal{D}^{i}_{k}\leftarrow$ GenerateData($\mathcal{D}_{E},\mathcal{D}^{-i}_{k-1}$)
+13       $\beta^{i}_{k}\leftarrow$ LearnBestResponse($\mathcal{D}^{i}_{k},\beta^{i}_{k-1}$)
+14       $\mathcal{D}^{i}_{k}\leftarrow$ UpdateAveragePolicy($\mathcal{D}^{i}_{k-1},\beta_{k}^{i}$)
+15        Update collection of BRs $\Pi_{i}\leftarrow\Pi_{i}\cup\{\beta_{k}^{i}\}$;
+16
+17      end for
+18
+19   end for
+20 return Aggregate($\Pi$)
+21
+22
+231exFunction GenerateData($\mathcal{D}_{E},\mathcal{D}^{-i}_{k-1}$):
+24   Get probability of $\pi_{k-1}^{-i}$ from $\mathcal{D}^{-i}_{k-1}$;
+25
+26  Calculate $w_{\pi_{k-1}^{-i}}$ following Equation 3;
+27
+28 $\mathcal{D}^{i}_{k}\leftarrow$ WeightData($\mathcal{D}_{E},w_{\pi_{k}^{-i}}$);
+29
+30 return $\mathcal{D}^{i}_{k}$
+31
+32
+331exFunction LearnBestResponse($\mathcal{D}^{i}_{k},\beta^{i}_{k-1}$):
+34   Initialize policy parameters with $\beta^{i}_{k}\leftarrow\beta^{i}_{k-1}$;
+35
+36  Optimize $\beta^{i}_{k}$ with offline RL and data $\mathcal{D}^{i}_{k}$;
+37
+38 return $\beta^{i}_{k}$
+39
+40
+411exFunction UpdateAveragePolicy($\mathcal{D}^{i}_{k-1},\beta^{i}_{k}$):
+42 for $(s,a,\pi_{k-1}^{i}(s,a))\in\mathcal{D}^{i}_{k-1}$ do
+43      Calculate $\pi_{k}^{i}(s,a)$ with $\beta^{i}_{k},\pi^{i}_{k-1}$ following Equation 1, and update dataset $\mathcal{D}^{i}_{k}$;
+44
+45   end for
+46 return $\mathcal{D}^{i}_{k}$
+47
+48
+Algorithm 2 Offline Fictitious Self-Play
+Appendix B Related Work
 In competitive games, also known as zero-sum games, people have developed a series of learning algorithms, where finding Nash equilibrium with self-play is the major learning paradigm for this problem.
-Fictitious self-play (FSP) (Heinrich et al., 2015) combines fictitious play (FP) (Brown, 1951) with self-play and provably converges to an NE in extensive-form games.
-Counterfactual regret minimization (CFR) (Bowling et al., 2015) combines regret minimization with self-play, first solving an imperfect-information game of real-world scale.
-DeepNash (Perolat et al., 2022) extends regularised Nash dynamics with RL and converges to an approximate NE in Stratego.
-Policy-Space Response Oracles (PSRO) (Lanctot et al., 2017) learn to find NE by iteratively learning best responses to its policies’ population, which can also be seen as a population-based self-play.
-Some works (Silver et al., 2017; Ye et al., 2020; Yang et al., 2022) of large-scale games aim for a higher winning percentage or returns and try to exploit opponents, not pursuing strict NE, still based on the self-play paradigm.
-In the offline learning paradigm, behavior cloning (BC) is the simplest form that directly imitates the sampling strategy of datasets.
-The performance of BC is limited by the sampling strategy in the dataset, while offline RL can surpass this limitation by maximizing returns.
+Fictitious self-play (FSP) (Heinrich, Lanctot, and Silver 2015) combines fictitious play (FP) (Brown 1951) with self-play and provably converges to an NE in extensive-form games.
+Counterfactual regret minimization (CFR) (Bowling et al.
+2015) combines regret minimization with self-play, first solving an imperfect-information game of real-world scale.
+DeepNash (Perolat et al.
+2022) extends regularised Nash dynamics with RL and converges to an approximate NE in Stratego.
+Policy-Space Response Oracles (PSRO) (Lanctot et al.
+2017) learn to find NE by iteratively learning best responses to its policies’ population, which can also be seen as a population-based self-play.
+Some works (Silver et al.
+2017; Ye et al.
+2020; Yang et al.
+2022) of large-scale games aim for a higher winning percentage or returns and try to exploit opponents, not pursuing strict NE, still based on the self-play paradigm.
+In the offline learning paradigm, behavior cloning (BC) is the simplest form that directly imitates the sampling policy of datasets.
+The performance of BC is limited by the sampling policy in the dataset, while offline RL can surpass this limitation by maximizing returns.
 Offline RL faces the challenge of the extrapolation error on out-of-distribution states and actions.
-Batch-constrained deep q-learning (BCQ) (Fujimoto et al., 2019b) and conservative q-learning (CQL) (Kumar et al., 2020) mitigate this issue by constraining the gap between policy and data distribution.
-Critic regularized regression (CRR) (Wang et al., 2020) solve this problem by weighted behavior cloning.
-In offline competitive scenarios, BC is a viable algorithm, still limited by the quality of sampling strategy in datasets.
+Batch-constrained deep q-learning (BCQ) (Fujimoto, Meger, and Precup 2019) and conservative q-learning (CQL) (Kumar et al.
+2020) mitigate this issue by constraining the gap between policy and data distribution.
+Critic regularized regression (CRR) (Wang et al.
+2020) solve this problem by weighted behavior cloning.
+In offline competitive scenarios, BC is a viable algorithm, still limited by the quality of sampling policy in datasets.
 Mathieu et al.
 (2023); Qu et al.
-(2023) directly learn strategies by single-agent offline RL without self-play.
-This way also requires high-quality behavior strategies in datasets and is easily overfitted to fixed opponents.
-Cui & Du (2022a); Zhong et al.
-(2022); Cui & Du (2022b) propose theoretically feasible algorithms.
+(2023) directly learn policies by single-agent offline RL without self-play.
+This way also requires high-quality behavior policies in datasets and is easily overfitted to fixed opponents.
+Cui and Du (2022b); Zhong et al.
+(2022); Cui and Du (2022a) propose theoretically feasible algorithms.
 Li et al.
@@
 Off-FSP implement a practical algorithm to find NE and has the potential to learn on partially covered datasets by integrating any variants of single-agent offline RL algorithms.
-3 Preliminaries
-3.1 Extensive-form Game
-Extensive-form games are a model of sequential interaction involving $n$ agents.
-Each player’s goal is to maximize his payoff in the game.
-At each step $t$ of extensive-form game, only one player observes his respective information states $s^{i}_{t}\in\mathcal{S}^{i}$ and suggests his action $a^{i}_{t}\in\mathcal{A}^{i}(s^{i}_{t})$.
-$\mathcal{S}^{i}$ is the set of information states of player $i$, and $\mathcal{A}^{i}(s)$ is the set of aviliable action at state $s$.
-The game’s randomization is defined by an extra player $c$, called chance.
-Chance follows a fixed randomized strategy that determines the distribution of the next state at its turns.
-The player function $P:\mathcal{S}\rightarrow\mathcal{N}\cup\{c\}$, with $\mathcal{N}=\{1,\dots,n\}$ denotes the set of players, determines who is to act at a given state.
-We assume the games with perfect recall.
-Each player’s current information state $s_{t}^{i}$ implies the trajectory sequence $(s_{1}^{i},a_{1}^{i},s_{2}^{i},a_{2}^{i},\dots,s_{t}^{i})$, including information states and actions led to $s_{t}^{i}$.
-In extensive-form games, each player plays following a behavioural strategy that maps all information states to distributions of available actions.
-The realization-plan (Von Stengel, 1996), $x(s_{t}^{i})=\Pi_{j=1}^{t-1}\pi^{i}(a_{j}^{i}|s_{j}^{i})$, describes the probability of reaching the information state $s_{t}^{i}$ following player $i$’s behavioural strategy, $\pi^{i}$.
-The strategy profile $\pi=(\pi^{1},\dots,\pi^{n})$, is a joint of all player’s behavioural strategy.
-$\pi^{-i}$ denotes the strategy profile of all players excepts $i$.
-The payoff of player $i$ is denoted by $R^{i}(\pi^{i},\pi^{-i})\in\mathbb{R}$, and $\sum_{i}R^{i}=0$ for zero-sum games.
-Given a fixed $\pi^{-i}$, best response $\textsc{BR}(\pi^{-i})$ is the strategy that has the highest payoff.
-A Nash equilibrium (NE) is a strategy profile that any strategy $\pi^{i}$ in this profile is a best response to the opponent’s profile $\pi^{-i}$.
-The $\epsilon$-best response ($\epsilon$-BR) and $\epsilon$-Nash equlibirium ($\epsilon$-NE) are approximations to above defination.
-$\epsilon$-BR is suboptimal by no more than $\epsilon$ compared with BR.
-Similarly, $\epsilon$-NE is a profile of $\epsilon$-BR.
-NashConv (Timbers et al., 2022), also called exploitability in two-player zero-sum games, evaluates the distance from $\pi$ to an NE, defined as $\sum_{i}R^{i}(\textsc{BR}(\pi^{-i}),\pi^{-i})-R^{i}(\pi^{i},\pi^{-i})$.
-Normal-form game is a simplified game description in game theory.
-Normal form represents the game by way of a matrix.
-An extensive-form game induces an equivalent normal-form game.
-In Sections 4.4 and 5.1, we discuss with toy examples of a normal-form game, rock-paper-scissors, and show the experimental results.
-Algorithm 1 Fictitious Self-Play
-Function FictitiousSelfPlay()
-      Initialize strategy $\pi_{0}$.  for $k=1$ to $K$ do
-           $\mathcal{D}_{k}\leftarrow$ GenerateData($\pi_{k-1}$) foreach player $i\in\mathcal{N}$ do
-                $\beta^{i}_{k}\leftarrow$ LearnBestResponse($\mathcal{D}_{k},\pi^{-i}_{k-1}$)
-          $\pi_{k}\leftarrow$ UpdateAverageStrategy($\pi^{i}_{k-1},\beta^{1}_{k},\dots,\beta^{n}_{k}$)
-     return average strategy $\pi_{K}$
-3.2 Fictitious Self-Play
-Fictitious Play (Brown, 1951) is a game-theoretic model for learning from self-play.
-Fictitious players iteratively compute the best responses to opponents’ average behavioural strategy and update their set of strategies.
-The average strategies of fictitious players are proven to converge to Nash equilibria in two-player zero-sum and many-player potential games.
-Generalised weakened fictitious play (Leslie & Collins, 2006) is an improvement of fictitious play that enables approximate best responses and perturbed average strategy updates.
-Generalised weakened fictitious play also provably converges to Nash equilibria in certain game classes and is suitable for machine learning framework.
-Original fictitious Play is defined in normal-form games, which requires exponential computations for extensive-form games.
-Heinrich et al.
-(2015) introduce Fictitious Self-Play (FSP) that extend fictitious play and generalised weakened fictitious play into the extensive-form game with linear time and space complexity.
-We briefly describe FSP at  Algorithm 1.
-In extensive-form games, the average strategies $\pi^{i}_{k}$ are updated by the realization-equivalence theorem.
-At the $k$-th iteration of FSP, the average behavioural strategy of player $i$ is
-$\begin{split}\forall s,a:~{}&\pi^{i}_{k}(a|s)=(1-\lambda)\pi^{i}_{k-1}(a|s)+\lambda\beta^{i}_{k}(a|s),\\ &\lambda=\frac{\alpha_{k}x_{\beta^{i}_{k}}(s)}{(1-\alpha_{k})x_{\pi^{i}_{k-1}}(s)+\alpha_{k}x_{\beta^{i}_{k}}(s)},\end{split}$ (1)
-where $\beta^{i}_{k}\in\epsilon_{k}\textsc{-BR}(\pi_{k-1}^{-i})$ is a best-response to opponent $\pi_{k-1}^{-i}$ and $\alpha_{k}$ is parameter to mix the best-response $\beta^{i}_{k}$ and previous average strategy $\pi^{i}_{k-1}$.
-Strategy $\pi^{i}_{k}$ is equivalent to choosing either strategy $\pi^{i}_{k-1}$ or $\beta^{i}_{k}$ before the beginning of each game, with probabilities $1-\alpha_{k}$ and $\alpha_{k}$ respectively.
-When $\alpha_{k}\rightarrow 0$ and $\epsilon_{k}\rightarrow 0$ as $k\rightarrow 0$, and $\sum_{k=1}^{\infty}\alpha_{t}=\infty$ for all $n$ players, the average strategy profile $\pi_{k}=(\pi^{1}_{k},\dots,\pi^{n}_{k})$ converge to an NE.
-In a common setting with $\alpha_{k}=\frac{1}{k}$, $\pi_{K}^{i}$ is equivalent to choosing all strategy $\beta^{i}_{1},\dots,\beta^{i}_{K}$ with the probability of $\frac{1}{K}$.
-The update of average behavioural strategy requires a traverse for all information state $s^{i}\in\mathcal{S}^{i}$, which is computationally expensive for large games and machine learning-based frameworks.
-FSP restored all the data of strategies $\beta^{i}_{1},\dots,\beta^{i}_{k}$ and trained an approximate average strategy with supervised learning.
-3.3 Offline Reinforcement Learning
-(a) Off-FSP and Off-SP.
-(b) Differences between FSP and Off-FSP.
-Figure 1: Illustration of Off-FSP, Off-SP and three essential functions. The green box in (a) is the Off-SP.
-Reinforcement learning (RL) agents aim to maximize the expected cumulative discounted reward in a Markov decision process (MDP).
-MDP is denoted as a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},\rho_{0},r,\gamma,T)$.
-$\mathcal{S},\mathcal{A}$ represent the state and action spaces.
-$\mathcal{T}(s_{t+1}|s_{t},a_{t}),r(s_{t},a_{t})$ represent the dynamics and reward function.
-$\rho_{0}$ is the distribution of initial state $s_{1}\sim\rho_{0}$.
-$\gamma\in[0,1]$ is the discount factor.
-Many reinforcement learning algorithms learn from transition tuples, $(s_{t},a_{t},r_{t},s_{t+1})$, where $s_{t},a_{t},r_{t}$ are the state, action and reward at time $t$, and $s_{t+1}$ is the next state after agent take action $a_{t}$ at state $s_{t}$.
-The expected cumulative discounted reward following policy $\pi$ can be formalized as the action-value function $Q(s_{t},a_{t})=\mathbb{E}_{\pi}[\sum^{\infty}_{i=t}\gamma^{i-t}r(s_{i},a_{i})]$.
-Model-based RL methods learn the dynamic models while model-free methods directly learn the Q function to state-action pairs.
-Given the fixed opponent $\pi_{-i}$ and an extensive form game, the game of player $i$ can be defined as a MDP $\mathcal{M}(\pi_{-i})$ (Silver & Veness, 2010; Greenwald et al., 2013).
-Player $i$’s information state defines the state, and his payoff function defines the reward.
-An $\epsilon$-optimal policy of the MDP, $\mathcal{M}(\pi^{-i})$, is also the $\epsilon$-BR to the strategy $\pi^{-i}$.
-Offline RL algorithm breaks the assumption that the agent can interact with the environment.
-The offline RL algorithm learns to maximize the expected cumulative reward based on a fixed dataset.
-The dataset is sampled following a behavior policy $\pi_{b}(a|s)$ in MDP $\mathcal{M}$.
-$\hat{\pi}_{b}(a|s):=\frac{\sum_{s,a\in\mathcal{D}^{i}}\mathbbm{1}[s=s,a=a]}{\sum_{s\in\mathcal{D}^{i}}\mathbbm{1}[s=s]}$ denote the empirical behavior policy, at all state $s\in\mathcal{D}^{i}$.
-In offline RL, the Q-function is trained only on actions sampled from offline datasets, which cannot cover all state and action spaces.
-Therefore, the Q-function may be erroneously overestimated at out-of-distribution (OOD) actions, which is called extrapolation error (Fujimoto et al., 2019b).
-Recent offline RL methods encourage the policy to learn on the support of training data or employ weighted behavior cloning to mitigate this error.
-4 Offline Fictitious Self-Play
-In this section, we introduce the Offline Self-Play (Off-SP) learning framework and give an implementation as Offline Fictitious Self-Play (Off-FSP) algorithm to find an $\epsilon$-Nash equilibrium with a fixed dataset.
-In the online scenario, self-play is a process in which a strategy iteratively maximizes cumulative rewards against itself, and fictitious self-play plays against the average of its past strategies.
-Corresponding to the concept of the online setting, Off-SP is a learning framework that enables strategy play against changing opponents with a fixed dataset.
-Off-FSP implement fictitious self-play on this learning framework with average strategies as the changing opponents and learn to find NE.
-Initially, we describe the offline datasets and make strong assumptions to simplify the problem.
-Based on the original FSP, as described in Algorithm 1, we derive the offline fictitious self-play by replacing three essential functions.
-In Section 4.2, we replace GenerateData by simulating play against different opponents with weighted datasets and replace LearnBestResponse with single-agent offline RL algorithm.
-The first two functions introduce Off-SP.
-In Section 4.3, we replace UpdateAverageStrategy by computing the average strategy based on samples and derive Off-FSP.
-Lastly, we discuss how offline RL helps Off-FSP to reduce the exploitability under real-world datasets.
-4.1 Assumptions of Offline Datasets
-Trajectory and Dataset.
-The trajectory of extensive-form games is a sequence of information states, actions, and player’s id.
-The trajectory also involves states and actions of chance $c$.
-The trajectory $\tau_{E}=\{s_{1},a_{1},p_{1},s_{2},\dots,s_{T},a_{T},p_{T}\mid p_{t}=P(s_{t})\},$
-We can assume the next information state for a prefix of trajectory and the payoff for a whole trajectory are definite.
-In other words, if we know the information sets of all players and chance, the game is deterministic.
-In player $i$’s perspective, the extensive-form game can be modeled as an MDP $\mathcal{M}$ given a fixed opponent $\pi^{-i}$ (Silver & Veness, 2010; Greenwald et al., 2013).
-In MDP $\mathcal{M}$, the dynamic function $\mathcal{T}(s^{i}_{t+1}\mid s^{i}_{t},a^{i}_{t})$ models the dynamics of opponents and the chance player.
-The observed MDP is non-deterministic because of imperfect information in a single player’s information state.
-To learn best-response with single-agent RL, we project the trajectory $\tau_{E}$ into player $i$’s perspective with a projection function $\tau^{i}=\mathcal{F}^{i}(\tau_{E})$.
-The projection $\mathcal{F}^{i}$ filters out the states corresponding to player $i$ while relabeling the time indices.
-$\begin{split}\tau^{i}&=\mathcal{F}^{i}(\tau_{E})=\{s_{t},a_{t}\mid\forall p_{t}\in\tau_{E},p_{t}=i\}\\ &=\{s_{1}^{i},a_{1}^{i},s_{2}^{i},\dots,s_{T^{\prime}}^{i},a_{T^{\prime}}^{i}\},\end{split}$
-where $T^{\prime}$ is the length of the player $i$’s trajectory $\tau^{i}_{\mathcal{M}}$.
-The subscripts of $s_{t}$ and $s_{t}^{i}$ are different and represent the time indices in the corresponding trajectories.
-For state $s^{i}_{t}\in\tau^{i}_{\mathcal{M}}$, we use function $I(s^{i}_{t})$ to denote the subscript of state in $\tau_{E}$.
-We denote opponent $j$’s latest state at player $i$’s state $s_{t}^{i}$ with function $\tau_{<}^{j}(s_{t}^{i})$.
-$\tau_{<}^{j}(s_{t}^{i})=s_{k},~{}\text{where}~{}{k=\max\{k^{\prime}<I(s^{i}_{t})\mid p_{k^{\prime}}=j\}}.$
-In order to offer a clear explanation of the stated components, we present an example trajectory in Figure 2.
-Figure 2: An example of extensive-form games and datasets.
-(Left) An example trajectory in extensive form game; (Right) Illustration of trajectories and dataset in extensive-form game and single-agent perspective.
-The orange area represents the player function.
-Blue parts represent the trajectories.
-The purple area represents the projection relationship under function $\mathcal{F}^{1}$.
-The green parts explain how function $\tau^{2}_{<}$ works.
-Based on the description of trajectory in both extensive-form games and single-agent perspective, the datasets consist of multiple corresponding trajectories.
-For extensive-form game, the dataset is $\mathcal{D}_{E}=\{\tau_{E}\}$.
-The dataset for player $i$ is $\mathcal{D}^{i}=\{\tau^{i}_{\mathcal{M}}=\mathcal{F}^{i}(\tau_{E})\mid\tau_{E}\in\mathcal{D}_{E}\}$.
-The $i$-th player’s dataset, derived from dataset $\mathcal{D}_{E}$, is denoted as $\mathcal{D}^{i}=\mathcal{F}^{i}(\mathcal{D}_{E})$.
-For briefly, we ignore the reward transformation in trajectories.
-Assumptions for Offline Learning.
-To introduce our method, we first assume the offline dataset $\mathcal{D}_{E}$ is both a fully covered dataset and a real-equivalence dataset.
-This assumption, crucial for the theoretical derivation of our methods, is not a prerequisite for practical application.
-The adaptability of our method to real-world datasets, unconstrained by these assumptions, is discussed in Section 4.4.
-The definitions of these two concepts are provided as follows.
-Definition 4.1 (Fully Covered Dataset).
-[…]
-When a dataset is fully covered, the trained learning algorithm will not suffer from OOD problems.
-Definition 4.2 (Real-Equivalence Dataset).
-For a dataset of extensive-form game $\mathcal{D}_{E}$ sampled following a behavior strategy $\pi_{b}$, dataset $\mathcal{D}_{E}$ is an real-equivalence dataset if $Pr(\tau_{E})=\frac{\sum_{\tau\in\mathcal{D}_{E}}\mathbbm{1}[\tau=\tau_{E}]}{|\mathcal{D}_{E}|},\forall\tau_{E}\in\mathcal{D}_{E},$ where $Pr(\tau_{E})$ is the probability of sampling trajectory $\tau_{E}$ with $\pi_{b}$ online.
-When a dataset is a real-equivalence dataset , sampling trajectories from the offline dataset $\mathcal{D}_{E}$ is equivalent to sampling from the extensive-form game with strategy $\pi_{b}$ online and the right-hand side of the equation is the probability density of $\tau_{E}$ in $\mathcal{D}_{E}$, denoted by $\mathcal{D}_{E}(\tau_{E})$.
-4.2 Learning Open-ended Best-Response
-Self-play and fictitious self-play iteratively interact with the game to generate data and utilize the data to learn best responses.
-In $k$-th iteration, the opponent for player $i$ is changed to $\pi^{-i}_{k-1}$, therefore, in order to learn the best response $\beta^{i}_{k}=\text{BR}(\pi^{-i}_{k-1})$, player $i$ must interact with $\pi^{-i}_{k-1}$ to generate new data $\mathcal{D}^{i}_{k}$.
-The process, from the perspective of player $i$, is equivalent to interacting with a new MDP $\mathcal{M}(\pi^{-i}_{k})$ and maximizing returns with RL to learn the corresponding best response (Greenwald et al., 2013).
-In offline settings, we are limited to using a fixed dataset sampled by behavioural policy $\pi_{b}$, in which RL methods can only obtain the best response $\text{BR}(\pi_{b}^{-i})$ for player $i$.
-In order to execute the fictitious self-play, we should generate player $i$ datasets under different MDP $\mathcal{M}(\pi^{-i}_{k})$.
-With importance sampling (Nachum et al., 2019), we can emulate sampling from another dataset $\mathcal{D}^{i}_{w}$ with weighting $w(d_{t})=\frac{\mathcal{D}^{i}_{w}(d_{t})}{\mathcal{D}^{i}(d_{t})}$ (Hong et al., 2023) on the original dataset $\mathcal{D}^{i}$, where $\mathcal{D}^{i}_{w}(d_{t})$ and $\mathcal{D}^{i}(d_{t})$ denote the probability density of tuple $d_{t}=(s^{i}_{t},a^{i}_{t},s^{i}_{t+1})$ in according datasets.
-This formulation gives the following equivalence:
-$\mathbb{E}_{d_{t}\sim\mathcal{D}^{i}_{w}}[L(d_{t};\theta)]\Leftrightarrow\mathbb{E}_{d_{t}\sim\mathcal{D}^{i}}[w(d_{t})L(d_{t};\theta)],$ (2)
-where $L(d_{t};\theta)$ is the loss function of an off-policy RL method, and $\theta$ is the parameter of RL policy to be optimized.
-With the assumption that dataset $\mathcal{D}_{E}$ is both a fully covered dataset and a real-equivalence dataset, sampling from it with importance sampling is equivalent to sampling from the online game with importance sampling.
-Theorem 4.3.
-For player $i$, the weight of transferring opponent from $\pi^{-i}_{b}$ to $\pi^{-i}$ is:
-$w(d_{t})=\frac{x^{-i}_{\pi^{-i}}(s_{j}){\pi^{-i}}(a_{j}|s_{j})}{x^{-i}_{\pi_{b}^{-i}}(s_{j}){\pi_{b}^{-i}}(a_{j}|s_{j})},~{}s_{j}=\tau_{<}^{-i}(s_{t+1}^{i}).$ (3)
-In order to simplify the phrase, we denote the strategies of $n-1$ opponents as an ensemble strategy $\pi^{-i}$ and refer to them as player $-i$.
-The proof of Theorem 4.3 is provided in Appendix A.
-Given an offline dataset $\mathcal{D}_{E}$, the behavior strategy $\pi_{b}$ can be approximated by an empirical behavioural policy $\tilde{\pi}_{b}$.
-In extensive-form games, we can estimate $\tilde{\pi}_{b}$ by counting or supervised learning policy.
-When considering a particular opponent, the weight $w(d_{t})$ assigned to a sample $d_{t}$ may shift to zero or a large value.
-As a result, when training RL with batched samples, the RL loss suffers from a large variance (Munos et al., 2016), leading to unstable training.
-To address this issue, we employ $\frac{w(d_{t})}{\sum_{d\in\mathcal{D}}w(d)}$ as the sampling probability rather than multiply the loss function by $w(d_{t})$ as shown in Equation 2.
-In this way, re-sampling from fully covered real-equivalence datasets is identical to sample data in online games.
-With batches of data sampled from the dataset, we can apply an offline RL to learn best responses.
-Figure 0(b) and Figure 3 illustrates these two functions and Off-SP.
-In the function GenerateDataset, we first calculate $w(d)$ for all the samples $d_{t}$ and get a re-weighted dataset $\mathcal{D}_{k}^{i}$ for each player $i$.
-In the function LearnBestResponse, offline RL algorithms repeatedly sample a batch of data and optimize the learned policy for $M$ times.
-These two functions derive the Off-SP.
-At each step $t$, Off-SP first generates dataset play against the current opponents and updates strategies towards the best response of the opponents.
-Figure 3: Offline Self-Play learning framework.
-4.3 Aggregate Average Strategy in Samples
-Before diving into Off-FSP, it is essential to review the mechanics of fictitious play in an online scenario.
-In norm-form games, fictitious play updates the new average strategy $\pi_{k+1}$ following Equation 1 after learning the best responses.
-To extend FP on extensive-form games, FSP gathers the trajectories of $\beta^{i}_{k}$ during each iteration and employs supervised learning to approximate $\pi^{i}_{k+1}$ for each player.
-However, Off-FSP does not require the current average strategy $\pi_{k}$ to sample new data, but only requires probability $\pi_{k}(s,a)$ for all state-action pairs in a fixed dataset.
-Therefore, an interactable average strategy $\pi_{k}$ is not necessary.
-We can just maintain the probability $\pi_{k}(s,a)$ in the datasets for the weighting technique.
-Traverse along the trajectory $\tau_{i}$, Off-FSP computes the probability $\pi^{i}_{k}(s,a)$ for player $i$ following Equation 1.
-Figure 0(b) shows one step of the function UpdateAverageStrategy.
-To facilitate the calculation, we save the probability into the current dataset $\mathcal{D}^{i}_{k}$, changing the sample into $d^{\prime}=(s,a,\pi^{i}_{k}(s,a))$.
-Although the weighting technique does not require a real instance of average behavioural strategy, Off-FSP requires an interactable strategy as output.
-One way is utilizing supervised learning to imitate the behavioural strategy with the weighting technique.
-When the dataset is not a fully covered dataset, this restricts strategies to exclusively choosing actions that are present in the dataset.
-In our experiments, we opted for an alternative method that involves aggregating all policies within a collection $\Pi$.
-Before the beginning of each game, one of the strategies in $\Pi$ is chosen with a specific probability and the payoff of the aggregation is the expectative payoff over all possible strategies.
-This approach maintains the option of picking actions that are out of the distribution but necessitates a larger storage capacity for storing all strategies.
-Algorithm 2 and Figure 0(a) present the pipeline of Offline Fictitious Self-play.
-Algorithm 2 Offline Fictitious Self-Play
-Function Off-FictitiousSelfPlay($\mathcal{D}_{E}$):
-      Estimate the empirical behavioural policy $\tilde{\pi}_{b}$$\Pi_{i}\leftarrow\{\pi_{b}^{i}\},\quad\pi_{0}\leftarrow\tilde{\pi}_{b}$$\mathcal{D}^{i}_{0}\leftarrow\{(s,a,\pi^{i}_{0}(s,a))|(s,a)\in\mathcal{F}^{i}(\mathcal{D}_{E})\}$for $k=1$ to $K$ do
-           for each player $i\in\mathcal{N}$ do
-                $\mathcal{D}^{i}_{k}\leftarrow$ GenerateDataset($\mathcal{D}_{E},\mathcal{D}^{-i}_{k-1}$) $\beta^{i}_{k}\leftarrow$ LearnBestResponse($\mathcal{D}^{i}_{k},\beta^{i}_{k-1}$) $\mathcal{D}^{i}_{k}\leftarrow$ UpdateAverageStrategy($\mathcal{D}^{i}_{k-1},\beta_{k}^{i}$) Update collection of BRs $\Pi_{i}\leftarrow\Pi_{i}\cup\{\beta_{k}^{i}\}$
-     return Aggregate($\Pi$)
-Function GenerateDataset($\mathcal{D}_{E},\mathcal{D}^{-i}_{k-1}$):
-[…]
-Function LearnBestResponse($\mathcal{D}^{i}_{k},\beta^{i}_{k-1}$):
-      Initialize policy parameters with $\beta^{i}_{k}\leftarrow\beta^{i}_{k-1}$  Optimize $\beta^{i}_{k}$ with offline RL and data $\mathcal{D}^{i}_{k}$return $\beta^{i}_{k}$
-Function UpdateAverageStrategy($\mathcal{D}^{i}_{k-1},\beta^{i}_{k}$):
-      for $(s,a,\pi_{k-1}^{i}(s,a))\in\mathcal{D}^{i}_{k-1}$ do
-           Calculate $\pi_{k}^{i}(s,a)$ with $\beta^{i}_{k},\pi^{i}_{k-1}$ following Equation 1, and update dataset $\mathcal{D}^{i}_{k}$
-     return $\mathcal{D}^{i}_{k}$
-4.4 Learning in Real-world Datasets
-Considering an offline dataset sampled from extensive-form games, the two assumptions in Section 4.1 cannot be completely fulfilled.
-Initially, a discrepancy in the probability of trajectories is invariably present between game and sampled datasets.
-This results in an irremovable error in learning best responses.
-Leslie & Collins (2006) proposed weakened fictitious play that allows suboptimal $\epsilon$-best responses in the learning process.
-The discrepancy does not totally make fictitious play fail but affects the distance of the final solution from equilibrium.
-In practice, sampling a large dataset can mitigate the issue.
-The sample complexity is a common problem discussed in previous works about offline learning (Cui & Du, 2022a; Yan et al., 2022).
-Secondly, it is not realistic to gather a fully covered dataset, especially in large-scale games.
-Previous works have discussed the requirements of the coverage to find NE (Cui & Du, 2022a; Li et al., 2022).
-However, their assumption is too restrictive to be fulfilled in real-world problems.
-In contrast, this work discusses how to minimize the distance to NE, i.e.
-the exploitability, under a given dataset, which is a more realistic setting.
-In fictitious self-play, a straightforward way to reduce exploitability is to learn better $\epsilon$-best responses at each step, which is the same for Off-FSP.
-The LearnBestResponse function can utilize any variants of the offline RL algorithms, which effectively address the issue of incomplete data set coverage.
-Figure 4: Example partially covered datasets for Rock-Paper-Scissors.
-Numbers in the grids show the probability density of different samples.
-The red dashed boxes indicate the probability of different actions for corresponding behavioural strategies.
-(Left) A partially covered example with infinite samples.
-(Right) A partially covered example with finite samples.
-Considering infinite samples, the missing of some $(s,a)$ is caused by $\pi_{b}(s,a)=0$ where $\pi_{b}$ is the behavior strategy to sample the dataset.
-We show this case in a simple matrix game, Rock-Paper-Scissors(RPS), with the first example of Figure 4.
-In this case, the covered part of the matrix can be seen as a subgame, where agents can only choose actions from observed ones.
-The payoff of the subgame is known.
-A conservative solution for this dataset is to find the Nash equilibrium of the subgame.
-However, if the lower bound of payoff in this game is known, allowing the selection of out-of-distribution actions may be a more reasonable setting.
-With the knowledge that payoffs of RPS are in the range of $[-1,1]$, player $1$ consistently selects paper as it yields a payoff of $1$.
-For player $2$, the only action, rock, in the dataset gets a payoff of $-1$, and choosing OOD actions may be a better selection because the rock is already the worst action.
-Some offline RL algorithms, such as CQL, give the OOD actions low rewards rather than set hard restrictions on these actions, which enables the learned policy to choose OOD actions.
-The experimental results in  Sections 5.1 and 5.2 show that restricting the policy space to subgames leads to higher exploitability.
-[…]
-5 Experiments
-In this section, we design a series of experiments to show the performance of Off-FSP and investigate the OOD issue in offline competitive games.
-We first evaluate Off-FSP on toy examples in a matrix game, rock-paper-scissors.
-We compared Off-FSP against the state-of-the-art offline RL methods for competitive games, including OEF (Li et al., 2022), Behavior Cloning (BC), and other single-agent Offline RL methods, on an imperfect-information poker game, Leduc Hold’em.
-We also conduct ablation studies on Off-FSP in  Section D.2.
-To demonstrate the compatibility of our approach with existing single-agent offline RL algorithms, we combine Off-FSP with three methods: CQL (Kumar et al., 2020), BCQ (Fujimoto et al., 2019b) and CRR (Wang et al., 2020).
-For variants with CQL, denoted as CQL-$\alpha$, we conduct different hyperparameters $\alpha$ which indicates the penalty level of choosing OOD actions.
-Higher $\alpha$ means less OOD actions performed by the policy.
-When $\alpha=0$, there is no penalty, and CQL degrades to DQN.
-More details about the experiments are provided in Appendix C.
-5.1 Rock, Paper, Scissors
-To provide empirical evidence for our statements in Section 4.4 and show the effectiveness of our methods, we first give results on three datasets of rock-paper-scissors.
-The first dataset, FULL, is randomly sampled from a uniform behavioural strategy, $\pi_{b}=(\frac{1}{3},\frac{1}{3},\frac{1}{3})$, which is a fully covered dataset and an approximation of ideal dataset.
-P1 and P2 are partially covered datasets described in Figure 4.
-Figure 5 shows the NashConv within $500$ iterations and the average strategies of player $2$ after the end of training.
-In FULL dataset, both CQL-00 and CQL-$1$ with Off-FSP are converged to a $\epsilon$-NE with a small gap with online fictitious play.
-In the partially covered P2 dataset, only CQL with $\alpha>0$ can properly address the OOD issues successfully converges to a $\epsilon$-NE, while CQL-00 is misleading by one single sample of Rock2 and selects Rock2 with a probability of $1$.
-In the P1 scenario, employing CQL with a smaller $\alpha$ value results in a reduced NashConv.
-This outcome is attributed to the allowance of OOD actions, as elaborated in Section 4.4.
-The outcomes from experiments P1 and P2 demonstrate the complexities inherent in OOD problems.
-The findings suggest that neither arbitrary selection nor total exclusion of OOD actions is optimal, emphasizing the need for sophisticated offline algorithms.
-Figure 5: Results of RPS. (Left) NashConv. (Right) Player $2$’s Average Strategy.
-5.2 Leduc Hold’em
-(a) NashConv
-(b) OOD ratio
-(c) NashConv w/o OOD actions
-(d) Payoff
-Figure 6: Results of Leduc Hold’em
-We now show the results compared with existing algorithms for offline competitive games.
-We collect 5 datasets on Leduc Hold’em with different quality.
-Mix and Random are nearly fully covered datasets, which are sampled by a random policy and aggregate strategy with expert and random policy.
-P20, P10, and P5 are partially covered datasets sampled from mixed strategies of $20,10,5$ best responses.
-These best responses are randomly picked from the training process of online PSRO.
-The more BR picked for sampling, the more state-action pairs are covered by the datasets.
-We visualize the ratio of covered state-action pair for all datasets in Appendix D.
-All these datasets contain 10 0001000010\,00010 000 trajectories.
-Mix and P20 datasets sampled from a high-quality strategy, so we can easily obtain a good strategy by supervised learning their empirical behavioral strategies $\tilde{\pi}_{b}$.
-This paradigm is also called behavior cloning, which is a popular baseline for offline learning.
-We evaluate state-of-the-art offline learning baselines and variants of Off-FSP on all the datasets.
-Figure 5(a) shows the NashConv of all the algorithms, and all variants of our method outperform other baselines.
-The result also demonstrates the ability of Off-FSP to robustly integrate with any existing offline RL methods.
-BC supervised learning the empirical behavioral strategies $\tilde{\pi}_{b}$ of all datasets, which shows Off-FSP can also outperform the sampling strategy.
-Mix and P20 datasets are both sampled from high-quality strategies, while the Random dataset is sampled from a random strategy.
-OEF-DCFR and OEF-PSRO are two variants of OEF (Li et al., 2022), which is a model-based offline algorithm for finding NE.
-OEF mixed its strategy with BC, evaluated it in online games multiple times and used the minimum NashConv as its results.
-This operation involves interaction with online games, which violates the offline scenario.
-To fair comparison, we remove the mixing operation in our main experiment.
-The results with mixing are shown in  Appendix D.
-OEF operates based on the assumption of full data coverage, so OEF strategies have a high probability of performing OOD actions on partially covered datasets, as shown in Figure 5(b).
-To investigate the influence of OOD actions, we restrict all the strategies on the observed state-action pairs and evaluate these restricted strategies.
-Figure 5(c) demonstrates significant performance degradation in some methods.
-We also learned an online CFR to find a $\epsilon$-NE on the subgames of different datasets.
-For partially covered datasets, the $\epsilon$-NE shows large NashConv values.
-This indicates that directly solving NE in subgames is insufficient for resolving the offline learning problem in competitive games.
-In previous works (Mathieu et al., 2023; Qu et al., 2023), directly applying single-agent Offline RL on the fixed dataset is also an attempt to solve offline competitive games.
-This paradigm is identical to learning best responses to the behavioural strategies $\pi_{b}$.
-These strategies overfit their opponents and are vulnerable to exploitation.
-These works only pay attention to strategies with higher returns and do not discuss NE.
-Therefore, we show the payoff of variants of Off-FSP playing with corresponding single-agent Offline RL algorithms, as Figure 5(d).
-Our methods show positive payoff to all the single-agent methods.
-6 Conclusion
-In this paper, we study offline multi-agent reinforcement learning for competitive games and propose Off-SP and Off-FSP to enable single-agent offline RL algorithms to be applied in this scenario.
-We find that Off-FSP can approximate NE even with partially covered datasets.
-Extensive experiments show that all variants of Off-FSP significantly outperform state-of-the-art baselines in different datasets of an imperfect-information two-player zero-sum game, Leduc Hold’em.
-Impact Statements
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix A Derivation of Theorem
+Appendix C Derivation of Theorem
 Proof.
 Now we derive the weight when changing the opponent to $\pi_{-i}$.
-Since we do not need to adjust player $i$’s strategy, player $i$’s strategy is still considered as $\tilde{\pi}_{b}^{i}$.
+Since we do not need to adjust player $i$’s policy, player $i$’s policy is still considered as $\tilde{\pi}_{b}^{i}$.
 In player $i$’s perspective, the reach probability of a tuple $d_{t}=(s^{i}_{t},a^{i}_{t},s^{i}_{t+1})$ is
-$Pr(d_{t}|\pi^{-i})=\Pi_{j=0}^{t}\pi^{i}(s^{i}_{j},a^{i}_{j})\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})\cdot\pi^{i}(a^{i}_{t+1}|s^{i}_{t+1}),$
+$Pr(d_{t}|\pi^{-i})=Pr(s^{i}_{0})\pi^{i}(a^{i}_{t+1}|s^{i}_{t+1})\cdot\Pi_{j=0}^{t}\pi^{i}(s^{i}_{j},a^{i}_{j})\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j}),$
 where $\mathcal{T}_{\pi^{-i}}$ is the transition function under MDP $\mathcal{M}(\pi^{-i})$. So we have:
 $w_{\pi^{-i}}(d_{t})$ $=\frac{Pr(d_{t}|\pi^{-i})}{Pr(d_{t}|\tilde{\pi}^{-i}_{b})}$
-$=\frac{\Pi_{j=0}^{t}\tilde{\pi}_{b}^{i}(s^{i}_{j},a^{i}_{j})\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})}{\Pi_{j=0}^{t}\tilde{\pi}_{b}^{i}(s^{i}_{j},a^{i}_{j})\mathcal{T}_{\tilde{\pi}_{b}^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})}\cdot\frac{\tilde{\pi}_{b}^{i}(a^{i}_{t+1}|s^{i}_{t+1})}{\tilde{\pi}_{b}^{i}(a^{i}_{t+1}|s^{i}_{t+1})}$
+$=\frac{Pr(s^{i}_{0})\tilde{\pi}_{b}^{i}(a^{i}_{t+1}|s^{i}_{t+1})}{Pr(s^{i}_{0})\tilde{\pi}_{b}^{i}(a^{i}_{t+1}|s^{i}_{t+1})}\cdot\frac{\Pi_{j=0}^{t}\tilde{\pi}_{b}^{i}(s^{i}_{j},a^{i}_{j})\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})}{\Pi_{j=0}^{t}\tilde{\pi}_{b}^{i}(s^{i}_{j},a^{i}_{j})\mathcal{T}_{\tilde{\pi}_{b}^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})}$
 $=\Pi_{j=0}^{t}\frac{\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})}{\mathcal{T}_{\tilde{\pi}_{b}^{-i}}(s^{i}_{j+1}|s_{j},a_{j})}$
-Considering $\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})$ and $\mathcal{T}_{\tilde{\pi}_{b}^{-i}}(s^{i}_{j+1}|s_{j},a_{j})$, both of these two items are the multiplication of a series of transition probabilities, of which only the probability of opponents’ strategies $\pi^{-i}$ is different.
+Considering $\mathcal{T}_{\pi^{-i}}(s^{i}_{j+1}|s^{i}_{j},a^{i}_{j})$ and $\mathcal{T}_{\tilde{\pi}_{b}^{-i}}(s^{i}_{j+1}|s_{j},a_{j})$, both of these two items are the multiplication of a series of transition probabilities, of which only the probability of opponents’ policies $\pi^{-i}$ is different.
 So the above equation can be simplified to:
@@
 where $s_{j}=\tau_{<}^{-i}(s_{t+1}^{i})$. ∎
-Appendix B Baselines
+Appendix D Baselines
 Behaviour Cloning (BC).
@@
 Conservative Q-Learning (CQL)
-Conservative Q-Learning (Kumar et al., 2020) learns a conservative Q-function such that the expected value of a policy under this Q-function lower-bounds its actual value in order to alleviate overestimation of values induced by the distributional shift between dataset and the learned policy.
+Conservative Q-Learning (Kumar et al.
+2020) learns a conservative Q-function such that the expected value of a policy under this Q-function lower-bounds its actual value in order to alleviate overestimation of values induced by the distributional shift between the dataset and the learned policy.
 In practice, we solve such an optimization problem to update our Q-function, where the first term is the penalty for lower-bounding the Q-function, and the second term is the Bellman error.
@@
 Critic Regularized Regression (CRR)
-Critic Regularized Regression (Wang et al., 2020) handles the problem of offline policy optimization by value-filtered regression.
+Critic Regularized Regression (Wang et al.
+2020) handles the problem of offline policy optimization by value-filtered regression.
 It selectively imitates the dataset by choosing a function $f$ that is monotonically increasing in $Q_{\theta}$.
-[…]
+$\underset{\pi}{\arg\max}\ \mathbb{E}_{(\mathbf{s},\mathbf{a})\sim\mathcal{D}}\left[f\left(Q_{\theta},\pi,\mathbf{s},\mathbf{a}\right)\log\pi(\mathbf{a}\mid\mathbf{s})\right]$ (5)
 Batch-Constrained deep Q-learning (BCQ)
-Batch-Constrained deep Q-learning (Fujimoto et al., 2019b) uses a generative model to learn the distribution of transitions in the dataset, samples several actions from it in the current state, perturbates them and chooses the one with highest Q value.
+Batch-Constrained deep Q-learning (Fujimoto, Meger, and Precup 2019) uses a generative model to learn the distribution of transitions in the dataset, samples several actions from it in the current state, perturbates them, and chooses the one with the highest Q value.
 In this way, they tried to minimize the distance of selected actions to the dataset and lead to states similar to those in the dataset.
-In the experiments, we use a discrete BCQ variant introduced in (Fujimoto et al., 2019a).
+In the experiments, we use a discrete BCQ variant introduced in (Fujimoto et al.
+2019).
 Single-Agent Offline RL
@@
 Offline Equilibrium Finding (OEF)
-Offline Equilibrium Finding (Li et al., 2022) is a model-based framework to find the game equilibrium in the offline setting.
+Offline Equilibrium Finding (Li et al.
+2022) is a model-based framework to find the game equilibrium in the offline setting.
 They directly train a model to describe the game dynamics and apply online equilibrium-finding algorithms like PSRO and DCFR (a deep learning version of CFR) to compute equilibrium.
 They also combine the behavior cloning policy with the model-based policy for improvement, where they mix these two policies with different weights and evaluate each weight in the online game to find the best combination.
-Appendix C Details of experiments
-C.1 Datasets
-C.1.1 Rock, Paper, Scissors
-FULL is randomly sampled from a uniform behavioral strategy, $\pi_{b}=(\frac{1}{3},\frac{1}{3},\frac{1}{3})$, i.e., every player randomly picks each action with the same probability.
-P1 and P2 are generated according to the exact proportion described in Figure 4, e.g., P1 comprises 500 episodes of Rock to Rock and 500 episodes of Paper to Rock.
-Every dataset is composed of 1000 trajectories.
-C.1.2 Leduc Hold’em
-Random is sampled by uniformly randomly picking action from all legal actions in every decision node.
-Mix is sampled from a mixed strategy, which is a 3-to-1 mix of expert and random strategies.
-The expert strategy is trained by PSRO with 30 iterations and is a mix of 30 deterministic policies according to a specific ratio.
-Partial dataset (5-10-20) is sampled by first picking 5, 10, and 20 policies from the 30 components of the expert strategy and then uniformly mixing them to form the sampling strategy.
-Every dataset comprises 10000 trajectories except those used in ablation study about the dataset size.
-C.2 Variants of Off-FSP
-Off-FSP-CQl, Off-FSP-CRR, and Off-FSP-CQL are different variants of our methods that integrate with different single-agent offline RL methods in the function of LearnBestResponse.
+Appendix E Details of experiments
+Datasets
+Datasets of Rock, Paper, Scissors
+D1 is randomly sampled from a policy, $\pi_{b}=(0.6,0.2,0.2)$. D2 is generated according to the exact proportion described in Figure˜3. Every dataset is composed of 1000 trajectories.
+Datasets of Extensive-form Games
+The Nash Conv of expert policies are $0.86,0.26,1.01$ in Leduc Poker, Large Kuhn Poker, and Oshi Zumo, separately.
+Every dataset comprises 10 000 trajectories The size of samples in datasets is far less than the number of samples required by online algorithms.
+Taking PSRO as an example, training BR requires about 10 000 trajectories at each iteration, and additional samples are needed to evaluate the new BR.
+Rock, Paper, Scissors
+As described in Figure˜3, the first dataset, D1, is randomly sampled from a uniform policy, $\pi_{b}=(0.6,0.2,0.2)$, which is a fully covered dataset and an approximation of the ideal dataset.
+D2 are partially covered datasets, which are constructed following the probability.
+Figure˜3 shows the NashConv within $500$ iterations and the average policies of player $2$ after the end of training.
+In D1 dataset, BC learns a suboptimal policy with an exploitability of $0.4$, while the exploitability of Off-FSP is less than $0.1$ after $500$ steps.
+Off-FSP-DQN is the ablation variant of Off-FSP without the offline RL algorithm, to learn the best response.
+It is easy to be misled by the OOD actions in D2.
+Off-FSP-CQL is the default setting of Off-FSP, which successfully learns a robust policy with low exploitability in both D1 and D2 datasets.
+Extensive-form Games
+Leduc Poker is a simplified version of two-player Texas holdem. Leduc Poker is a widely-used imperfect information normal-form game as a testbed for online algorithms.
+Large Kuhn Poker is a variant of Kuhn Poker where an initial pot for each player is $5$.
+Each player has four actions: Fold, Check, Call, and Raise $1$ pot.
+Only in the first $8$ steps can players raise.
+Oshi-Zumo is a deterministic board game in which two players repeatedly bid to push a token off the other side of the board (Buro 2004).
+The size of the board is $3$.
+The initial number of coins for each player is $4$.
+The maximum game horizon is $6$.
+Ball Defence Game
+The robot arm in the ball defence game is Agilex PiPER, a 6-DOF arm.
+It is equipped with an RGBD camera, Zed mini, to perceive the position of the blue ball.
+We use color segmentation with 1344x376 resolution and 100 FPS to detect the ball.
+The size of each basket, i.e., the goals, is 20x17x10 cm, and the size of the black racket is 16x13 cm.
+In data collection, we restricted the action space of the robotic arm’s end effector to the area above the basket and discretized it into a 10-dimensional discrete space.
+This increases the success rate of data collection and ensures that the robot’s movements will not cause any damage to the experimental setup.
+In the evaluation phase, we invited 10 human players to play the ball defence game and evaluate the performance of all methods.
+For fair comparison, each human player first has a 3-minute warmup to familiarize themselves with this game, and then plays with each method in a random order.
+For each method, we only retain the first two results that each person throws at each goal.
+The total number of evaluations for each goal and each method is 20.
+Variants of Off-FSP
+Off-FSP-CQL, Off-FSP-CRR, and Off-FSP-BCQ are different variants of our methods that integrate with different single-agent offline RL methods in the function of LearnBestResponse.
 We then introduce details about these Offline RL methods.
 In CQL, we use a QRDQN as the base algorithm and a simple MLP as the Q network.
-Hence, the network’s input is the observation vector, and the output has a dimension of the number of quantiles times the number of actions.
+Hence, the network’s input is the observation vector, and the output has a dimension of the number of quantiles, multiplied by the number of actions.
 The hidden layer of MLP contains the same number of neurons.
 We use ReLU as the activation function between every linear layer.
-In CRR, the actor and critic networks have the same structure, first a shared feature network and then a MLP head.
-The shared feature network is a MLP with two layers, and the size of the second layer is the number of actions.
+In CRR, the actor and critic networks have the same structure, first a shared feature network, and then an MLP head.
+The shared feature network is an MLP with two layers, and the size of the second layer is the number of actions.
 All activation functions are ReLU.
-CQL penalty term is also added to the CRR objective function for better performance.
-In BCQ, the policy network and the imitation network have the same structure, which is also first a shared feature network and then an MLP head, exactly the same as CRR.
-C.3 Hyperparameters
-In OEF, all settings are the same as Li et al. (2022).
-In RPS, we use Adam optimizer with a learning rate of 0.005 and betas (0.9, 0.999), and the network has only one layer, and the number of quantiles in QRDQN is 100.
-The target network is not used.
-In each iteration, the network is trained five epochs and is updated 100 times each epoch with a batch size of 1000.
-In Leduc, the hyperparamters are shown in Table 1.
-All networks are optimized by Adam Optimizer (Kingma & Ba, 2017).
-The hidden layer number of CRR and BCQ describes the whole network, including the feature network and MLP head.
-The target network is used.
-Table 1: Hyperparameters used in Leduc
+The CQL penalty term is also added to the CRR objective function for better performance.
+In BCQ, the policy network and the imitation network have the same structure. They are also first a shared feature network and then an MLP head, which is exactly the same as CRR.
+Table 2: Hyperparameters used in Leduc Poker
 Hyperparameter CQL CRR BCQ
@@
 Adam $(\beta_{1},\beta_{2})$ (0.9, 0.999) (0.9, 0.999) (0.9, 0.999)
-Training steps per iteration $M$ 1 000 1 000 1 000
+Training iterations 2 000 2 000 2 000
+Training steps per iteration 1 000 1 000 1 000
 Target Network update frequency 100 100 100
@@
 QRDQN number of quantiles 100
-CQL $\alpha$ 0.1 0.1
+CQL $\alpha$ (mix) 2.0 0.1
+CQL $\alpha$ (population) 0.5 0.1
 BCQ unlikely action threshold 0.1
@@
 CRR ratio bound 20
-Appendix D Supplement Results
-D.1 Analysis of Datasets
-Figure 7: State-Action Cover Ratio
-We visualize the state-action cover ratio of different datasets in Figure 7
-D.2 Ablations
-(a) Value of $\alpha$
-(b) Size of Data
-Figure 8: Results of ablations.
-[…]
-D.3 Learning Curves of Off-FSP
-OEF evaluates strategies’ performance by allowing strategies to mix with BC by different weights.
-However, the results are decided by the best-mixed strategies.
-This is equivalent to allowing strategies to be interactively evaluated online, which is contrary to the setting of offline scenarios.
-And we cannot tell whether this result is caused by the original strategy or BC.
-We still provide the final results under this evaluation method in Figure 9, and our method still outperforms OEF.
-Figure 9: NashConv with evaluation method in OEF
+Table 3: Hyperparameters used in Large Kuhn Poker, Oshi Zumo and ball Defence Game.
+Hyperparameter Large Kuhn Poker Oshi Zumo Ball Defence
+Offline RL algorithm CQL CQL CQL
+Hidden layer number 4 4 4
+Hidden layer size 256 256 256
+Learning rate 0.0001 0.0001 0.0001
+Adam $(\beta_{1},\beta_{2})$ (0.9, 0.999) (0.9, 0.999) (0.9, 0.999)
+Training iterations 1 000 1 000 200
+Training steps per iteration 1 000 1 000 50
+Target Network update frequency 100 100 10
+Batch size 1024 1024 128
+QRDQN number of quantiles 100 100 100
+CQL $\alpha$ 0.1 0.01 0.1
+Hyperparameters
+For OEF baseline, all settings are the same as Li et al. (2022).
+In RPS, we use Adam optimizer with a learning rate of 0.005 and betas (0.9, 0.999), and the network has only one layer, and the number of quantiles in QRDQN is 100.
+The target network is not used.
+In each iteration, the network is trained for five epochs and is updated 100 times each epoch with a batch size of 1000.
+In extensive-form games, the hyperparameters are shown in Tables˜2 and 3.
+All networks are optimized by Adam Optimizer (Kingma and Ba 2017).
+The number of hidden layers in CRR and BCQ describes the whole network, including the feature network and MLP head.
+The target network is used.
+Our experiments run on AMD EPYC 7302 CPU and 3080Ti GPUs.
+In the ball defence game, the hyperparameters are shown in Table˜3.
+To accelerate the training process, we use softmax with temperature 0.5 to calculate the probabilities of each action.
+In the final evaluation, we only keep a checkpoint of the robot policy at 100, 125, 150, 175, and 200 iterations and randomly pick one of them each time.
+Appendix F Explanation on RPS experiments
+As described in Figure˜3, the first dataset, D1, is randomly sampled from a uniform behavioural strategy, $\pi_{b}=(0.6,0.2,0.2)$, which is a fully covered dataset and an approximation of the ideal dataset.
+D2 are partially covered datasets, which are constructed following the probability.
+Figure˜3 shows the NashConv within $500$ iterations and the average strategies of player $2$ after the end of training.
+In D1 dataset, BC learns a suboptimal strategy with an exploitability of $0.4$, while the exploitability of Off-FSP is less than $0.1$ after $500$ steps.
+Off-FSP-DQN is the variant of Off-FSP by using DQN, an online RL algorithm, to learn the best response.
+The algorithm only works well in ideal datasets such as D1, and it is easy to be misled by the OOD actions in D2.
+Off-FSP-CQL is the default setting of Off-FSP, which successfully learns a robust strategy with low exploitability in both D1 and D2 datasets.
+Appendix G Supplement Results
+Learning Curves of Nash Conv for Extensive-form Games
+Figure 10: Learning Curves of our main experiments.
+Figure˜10 shows all the Nash Conv’s learning curves of the main experiments in Section˜5.1. CQL is Q-learning algorithms with Q function only.
+Figure 11: Results on Extensive-Form Games with evaluation method in OEF. (Top) NashConv on Mix Datasets; (Bottom) NashConv on Population Datasets.
+Experimental Results Under Rules of OEF
+OEF evaluates policies’ performance by allowing policies to mix with BC by different weights.
+However, the results are decided by the best-mixed policies.
+This is equivalent to allowing policies to be interactively evaluated online, which is contrary to the setting of offline scenarios.
+It is difficult to figure out whether the original policy or BC causes this result.
+We still provide the final results under this evaluation method in Figure˜11, and our method still outperforms OEF.